### v3

# Video LLMs for Temporal Reasoning in Long Videos

Abstract
This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos. At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues. In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps into time-sensitive local features. Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation. The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos. Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation. Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
Figure 1: Previous video large language models (a-c) usually are not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling operation (a, b) or query aggregation (c) for aggregating global semantic information. In contrast, our model (d) includes a time-aware clip encoder, which extracts time-aware fine-grained cues from short-term clips sampled from a long-term input video, and a bidirectional long short-term memory module, which captures long-range temporal dependencies across multiple clips. Our extracted time-aware and multi-level features are crucial for temporal reasoning in long videos.
1 Introduction
00footnotetext: indicates joint first author.
{fawad,umer,hamza,zeeshan,huy}@retrocausal.ai.
Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time. Several video understanding applications require the ability of temporal reasoning. Examples include dense video captioning [45, 59, 52], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [22, 53, 2, 25], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes. Considerable efforts [52, 26, 29, 25] have been invested in developing separate models for solving individual tasks. These models usually have different architectures. Therefore, it is favorable to design a unified model for handling various tasks.
The past few years have witnessed the impressive instruction following ability and remarkable comprehension and generation capability of large language models (LLMs) [1, 4, 40, 30, 41, 42], which have emerged as a universal agent for performing various tasks. Video LLMs [21, 23, 27, 28, 56, 15, 38] which incorporate video encoders with LLMs for video understanding tasks have been introduced. These methods often represent a video by a fixed number of visual tokens, leading to reduced performance with long videos, and encode video frames and their timestamps separately and hence struggle to perform temporal reasoning tasks. Recently, a few works [32, 13, 14, 35] have developed video LLMs with temporal reasoning abilities, e.g., TimeChat [35] proposes to vary the number of final video tokens based on the input video length and jointly encode video frames and their timestamps. The above methods usually treat the entire video as a single clip [32, 13, 14, 35] and aggregate video tokens via pooling operation [14] and query aggregation [35], struggling to capture fine-grained information, especially for long videos.
In this paper, we propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos. Fig. 1 shows the differences in model architectures of TemporalVLM and previous works. Our model includes two key architectural contributions for temporal reasoning in long videos. Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting time-aware local features from each clip. Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips. As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos. Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies. Our IndustryASM dataset comprises of $4803$ videos with an average video duration of $119$ seconds. It covers in total $52$ diverse industry assembly tasks and includes timestamp and action labels. We convert the labels into chat samples with manually written instructions. To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses. Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [35] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
In summary, our contributions include:
  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos. Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into time-aware local features. The local features are then fed to a BiLSTM module to extract global features. Therefore, our features not only are time-sensitive but also capture effective local and global information.
  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures. Our videos are recorded on factory floors with actions and timestamps labeled by industrial engineers for time and motion analyses and temporal action segmentation evaluation. We will make our dataset available publicly.
  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to integrate LSTMs into video LLMs.
2 Related Work
Large Language Models. Over the past few years, LLMs [1, 4, 40, 30, 41, 42] have revolutionized the field of natural language processing. Initial methods employ encoder-decoder architectures [16, 34], whereas decoder-only models, e.g., GPT [33], demonstrate impressive performance and generalization in recent years. In addition, models with billions of parameters, e.g., GPT-4 [1], have been developed to push the boundary further, while instruction tuning techniques [31, 46, 47] have been widely used to enable models to generate contextually relevant responses. Here, we focus on video LLMs, which take videos as additional inputs and explore the use of LLMs for video understanding tasks.
Video Large Language Models. Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [21, 23, 27, 28, 56, 15, 38] have been developed for video understanding tasks. They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [37]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [41]) for generating responses. They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video. For example, VideoChat [21] extracts frame features via a visual transformer [37] and employs a query transformer (Q-Former) [20] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [56]. The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks. To address these drawbacks, methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [35] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps. These methods often consider the video as a single clip [27, 28, 21, 56] and aggregate video tokens via pooling operation [27, 28] and query aggregation [21, 56], overlooking fine-grained details. In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip. Also, we introduce a BiLSTM module for aggregating global features over multiple clips. To our best knowledge, our work is the first to exploit LSTMs for video LLMs.
Video Temporal Reasoning. Temporal reasoning plays an important role in many video understanding tasks, e.g., dense video captioning [45, 59, 52], temporal video grounding [48, 26], video highlight detection [19, 29], and temporal action segmentation [22, 53, 2, 25]. Previous works [52, 26, 29, 25] often focus on designing separate models for tackling individual tasks, while LLM-based models [32, 13, 14, 35] capable of handling multiple tasks have emerged recently. Our TemporalVLM model belongs to the second group. Also, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion analyses.
Long Video Understanding. There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information. Several methods have been developed for long video understanding, including efficient architectures [6, 17], temporal pooling/aggregation [36, 50], and dynamic clip selection [9, 10]. For vision-language understanding tasks, methods based on temporal alignment [11, 3] have been introduced, while memory techniques [51, 57] are often used in dense video prediction tasks. Recently, Weng et al. [49] propose LongVLM for long video understanding. It divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips. It does not utilize timestamps, which are crucial for temporal reasoning. In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features. Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps. The time-aware clip encoder is then employed on each clip along with its timestamps for extracting time-aware local features, which capture local fine-grained cues within each clip. Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips. Lastly, the large language model takes video and text tokens as inputs and generates responses.
Below we describe our first contribution, TemporalVLM, a video LLM for effective temporal reasoning in long videos. Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features. We provide an overview of our TemporalVLM model in Fig. 2.
3.1 Time-Aware Clip Encoder
Video LLMs often consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat [35] and $N^{v}_{f}=100$ frames in VTimeLLM [13]. Moreover, prior works usually use query aggregation, e.g., TimeChat [35], and pooling operation, e.g., Video-ChatGPT [28], to aggregate video tokens. Thus, they often perform well with short-term videos but miss out important fine-grained information for long-term videos, leading to degrading performance.
We propose a time-aware clip encoder (i.e., blue block in Fig. 2) for extracting local fine-grained cues within a clip. Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip. Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip. The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding time-aware local features. We adopt the time-aware frame encoder and sliding video Q-Former of TimeChat [35] as our time-aware clip encoder. The time-aware frame encoder uses a pre-trained image encoder [39] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features. These frame features are then fed to a sliding video Q-Former [35] for temporal modeling, yielding time-aware clip features for each clip. Recent works, e.g., LongVLM [49], leverage clip sampler and clip encoder for long video understanding but their encoders are not time-aware and hence are not effective for temporal reasoning.
3.2 Bidirectional Long Short-Term Memory
Our time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip. However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
We introduce a BiLSTM module (i.e., green block in Fig. 2) for computing global features across multiple clips. We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video. We then pass the sequence of local features to a BiLSTM module for aggregating global features. Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
$\mathbf{h}^{f}_{t}=\text{LSTM}(\mathbf{h}^{f}_{t-1},\mathbf{c}_{t}),$ (1)
$\mathbf{h}^{b}_{t}=\text{LSTM}(\mathbf{h}^{b}_{t+1},\mathbf{c}_{t}),$ (2)
where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$. The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as:
$\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}].$ (3)
Please refer to the textbook [18] for a detailed description. Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions. As empirically observed in Sec. 5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [43]. Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos. Unlike the recent LongVLM model [49], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec. 5.
The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$. Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM. Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
3.3 Large Language Model
The LLM (i.e., yellow block in Fig. 2) takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users. Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model. Following TimeChat [35], we use the checkpoint of the open-source LLaMA-2 7B model [42] and perform instruction tuning only. We optimize the below loss function:
$\mathcal{L}=-\log P_{\mathbf{\psi}}(\mathbf{X}^{r}|\mathbf{X}^{v},\mathbf{X}^{q})$ (4)
$=-\sum^{L}_{i=1}\log P_{\mathbf{\psi}}(x_{i}|\mathbf{X}^{v},\mathbf{X}^{q},\mathbf{X}^{r}_{<i}).$ (5)
Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response. Also, we employ LoRA [12] for efficient fine-tuning.
4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
Figure 3: Example videos with different camera viewpoints, actors, backgrounds, and activities from our IndustryASM dataset.
We follow prior works [32, 13, 14, 35] to test TemporalVLM on dense video captioning, temporal video grounding, and video highlight detection. To further evaluate our model, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies. To our best knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers. Below we provide some details and annotation process.
4.1 Dataset Statistics
Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration. In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average. Example videos are shown in Fig. 3. Due to space limits, we provide additional details of IndustryASM in our supplementary material.
4.2 Annotation Process
Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments. The action labels contain the same step naming conventions as used on factory floors during manufacturing processes. To ensure quality, each video is annotated by two labelers, i.e., one labeler obtains the labels for the video, while another labeler checks the labels. In cases of conflicts, both labelers discuss to address them. Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of roughly 92% between labelers. Given the framewise annotations we first convert them to timestamps by using the frame rate of the video. We manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using timestamps and action labels. Example instructions and responses are shown in Fig. S1 of our supplementary material.
5 Experiments
Tasks and Datasets. Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation. For fair comparisons with TimeChat [35], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [54] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [35] and Valley [27] datasets. The combined dataset contains a total of 142K long videos with diverse activities. We follow TimeChat [35] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples. We evaluate on YouCook2 [58] for dense video captioning, Charades-STA [8] for temporal video grounding, QVHighlights [19] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
Implementation Details. We follow TimeChat [35] to use LLaMA-2 7B [42] as the LLM, and ViT-G/14 from EVA-CLIP [39] as the image encoder. We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [56] checkpoints to initialize the video Q-Former. The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer. The parameters of the BiLSTM module and projection layers are initialized randomly. The LLM is fine-tuned using LoRA [12] with rank 32. We use 6 clips and sample 96 frames from each clip. We fine-tune our model on a subset of TimeIT [35] and Valley [27] for 7 epochs on a single 8-A100 (80 GB) machine. The parameters of the LLM and image encoder remain frozen throughout the fine-tuning. Please see our supplementary material for more details.
Competing Methods. We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [21], Video-LLaMA [56], Valley [27], TimeChat [35], and LongVLM [49]. Since LongVLM [49] is not trained on TimeIT [35] and Valley [27] and it is a multi-model pipeline (TimeChat [35] and TemporalVLM are end-to-end models), we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM model which we use for evaluation. Nevertheless, we compare with the original LongVLM model [49] on general video understanding in our supplementary material.
Evaluation Metrics. We utilize different sets of metrics for evaluating temporal reasoning tasks:
  • Dense video captioning: The overall story-level evaluation is done through SODA_c [7]. The event localization performance is measured using F1 Score. CIDEr [44] is used to measure the quality of the generated captions.
  • Temporal video grounding: To gauge the effectiveness of temporal video grounding, we use R@1(IoU = $\alpha$), which is the percentage of predictions with an intersection over union greater than $\alpha$ compared to the ground truth.
  • Video highlight detection: We use mean average precision (mAP) and HIT@1 at IoU= 0.75 to evaluate the accuracy of the saliency scores for key moments in videos.
  • Temporal action segmentation: We report the framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance. We use the intersection over union ratio to find the overlapping thresholds.
5.1 Zero-Shot Results
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [58] QVHighlights [19] Charades-STA [8]
SODA_c CIDEr F1 mAP HIT@1 R@1(IoU=0.5) R@1(IOU=0.7)
Valley [27] 0.1 0.0 1.5 10.9 15.2 4.7 1.6
Video-LLaMA [56] 0.0 0.0 0.1 11.3 15.6 2.7 1.2
VideoChat-Embed [21] 0.2 0.6 3.4 13.1 18.1 3.2 1.4
LongVLM [49] 0.8 2.5 12.3 11.1 15.0 13.9 6.1
TimeChat [35] 1.2 3.4 12.6 14.5 23.9 27.9 12.3
TemporalVLM (Ours) 1.2 3.7 13.1 16.4 31.3 30.1 13.2
Table 1: Zero-short results. We evaluate all methods on YouCook2 [58] for dense video captioning, QVHighlights [19] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
We first evaluate all methods in the zero-shot setting and present the results in Tab. 1.
5.1.1 Dense Video Captioning Results
For dense video captioning, it can be seen from Tab. 1 that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score. For example, we outperform TimeChat [35] by +0.3 on CIDEr and +0.5 on F1 Score. This is likely because TimeChat [35] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information. In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization. Moreover, our approach obtains superior results than LongVLM [49] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
5.1.2 Video Highlight Detection Results
The video highlight detection results in Tab. 1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1. In particular, it surpasses TimeChat [35] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [49] by +5.3 at mAP and by +16.3 at HIT@1. These results suggest the effectiveness of our model for video highlight detection. By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query. Though LongVLM [49] extracts multi-level features, the lack of a time-aware encoder hinders their results.
5.1.3 Temporal Video Grounding Results
It is evident from Tab. 1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7). Specifically, it outperforms TimeChat [35] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [49] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively. The results validate the benefits of our time-aware and multi-level features for temporal video grounding. Our time-aware clip encoder extracts fine-grained cues for temporal reasoning within each clip, while our BiLSTM module captures long-range temporal relationships between multiple clips to detect the given event within the video effectively. The lack of a time-aware encoder reduces the performance of LongVLM [49], while TimeChat [35] suffers from a lack of fine-grained information.
5.2 Supervised Results
We now evaluate all methods in the supervised setting.
5.2.1 Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results
Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
YouCook2 [58] QVHighlights [19] Charades-STA [8]
SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
LongVLM [49] 2.3 8.1 16.9 16.0 22.5 27.2 11.9
TimeChat [35] 3.1 10.3 19.5 21.7 37.9 46.7 23.7
TemporalVLM (Ours) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
Table 2: Supervised results. We first fine-tune all models on TimeIT [35] and Valley [27], and then perform task-specific fine-tuning and evaluation on YouCook2 [58] for dense video captioning, QVHighlights [19] for video highlight detection, and Charades-STA [8] for temporal video grounding. Best results are in bold, while second best ones are underlined.
We present the results of all methods in the supervised setting in Tab. 2. It can be seen from Tab. 2 that our TemporalVLM model achieves the best results across all tasks and metrics. As compared to the zero-shot results in Tab. 1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task. However, we acknowledge that task-specific models such as MMN [48] and QD-DETR [29] may outperform generalist models such as TimeChat [35] and our TemporalVLM model due to their task-specific design paradigms such as using a mutually matching network to directly model the relationship between the query text and video content and introducing specialized tokens and loss functions during fine-tuning.
5.2.2 Temporal Action Segmentation Results
Model F1@{10,25,50} Edit Acc
TimeChat [35] {11.1, 8.5, 4.4} 32.7 46.0
LongVLM [49] {17.1, 13.3, 7.3} 47.3 47.6
TemporalVLM (Ours) {22.3, 18.3, 11.1} 56.2 52.9
Table 3: Temporal action segmentation results in the supervised setting on IndustryASM. All models are fine-tuned on IndustryASM before evaluation. Best results are in bold, while second best ones are underlined.
We further test all methods on our IndustryASM dataset for temporal action segmentation. Since the activities in IndustryASM significantly differ from those in TimeIT [35] and Valley [27], we fine-tune all models on IndustryASM before evaluation. Tab. 3 presents the results. It is evident from Tab. 3 that TemporalVLM performs the best across all metrics. For example, it outperforms TimeChat [35] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [49] by +8.9 and +5.3 on Edit and Acc respectively. The results confirm the benefits of our time-aware and multi-level features for temporal action segmentation.
Multi-Clip Encoder vs. Time-Aware Encoder. Multi-clip encoders in LongVLM [49] and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat [35] on temporal action segmentation in Tab. 3. In contrast, time-aware encoders in TimeChat [35] and TemporalVLM are important to event localization and temporal modeling, resulting in better performance than LongVLM [49] on dense video captioning, video highlight detection, and temporal video grounding in Tabs. 1 and 2. Our TemporalVLM encoder is both multi-clip and time-aware, yielding the best performance across all tasks.
5.3 Ablation Results
# Clips SODA_c CIDEr F1
2 2.6 9.2 18.5
4 3.1 11.3 18.9
6 3.4 13.2 20.0
Table 4: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
Model # Layers SODA_c CIDEr F1
Average Pooling 1 1.7 6.0 16.1
Linear Layer 1 2.4 9.5 17.8
Transformer 2 2.3 6.8 15.6
LSTM 1 2.8 9.4 18.2
BiLSTM 2 3.4 13.2 20.0
Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
Below we study the effects of the number of short-term clips and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [58].
5.3.1 Impacts of the Number of Short-Term Clips
Tab. 4 presents the effects of the number of short-term clips on our TemporalVLM model. Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6. It is evident from Tab. 4 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
5.3.2 Impacts of the BiLSTM Module
To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab. 5. From Tab. 5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [28, 49], and traditional LSTM, which only performs the forward pass and hence relies only on past cues. In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues. Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [56, 35] alone, leading to worse results. This highlights the importance of BiLSTM even when video Q-Formers [56, 35] are being used. Finally, BiLSTM outperforms transformer [43], which we will discuss below.
(Bi)LSTM vs. Transformer. The permutation-invariant nature of the self-attention mechanism in transformer [43] leads to loss in temporal information of the input sequence. Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs [55]. In contrast, (Bi)LSTM processes data one by one and preserves the order of the sequence, making them better suited for temporal modeling. Similar to [55], we also observe in Tab. 5 that linear layer outperforms transformer [43].
5.4 Qualitative Results
Figure 4: Example dense video captioning results in the zero-shot setting on a YouCook2 [58] video. Inaccuracies are shown in red.
Figure 5: Example temporal action segmentation results in the supervised setting on an IndustryASM video. Colored segments represent predicted actions with a particular color denoting the same action across all models. Black represents background frames.
In this section, we discuss some qualitative results.
5.4.1 Dense Video Captioning Results
Fig. 4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [58] video. TimeChat [35] fails to predict any correct timestamps and tends to hallucinate. LongVLM [49] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions. In contrast, TemporalVLM produces considerably more accurate timestamps and captions. Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long video understanding.
5.4.2 Temporal Action Segmentation Results
Fig. 5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video. TimeChat [35] falls short in segmenting the video correctly and hallucinates actions not present in the video. LongVLM [49] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video. In contrast, TemporalVLM predicts action segments that are significantly closer to the ground truth.
6 Conclusion
We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos. Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features. The extracted features not only are time-sensitive but also contain both local and global cues. In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes. Our dataset is recorded on factory floors and includes action and timestamp labels provided by industrial engineers for time and motion studies and temporal action segmentation evaluation. Lastly, we conduct extensive experiments on TimeIT [35] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks. To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
Appendix A Supplementary Material
In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs. A.1 and A.4 respectively. Next, Sec. A.5 compares the sizes of our model, TimeChat [35], and LongVLM [49], while Sec. A.6 evaluates our model for general video understanding. We then present several qualitative results in Sec. A.7, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results. Finally, we discuss the limitations and societal impacts of our work in Sec. A.8.
A.1 IndustryASM Details
A.2 Dataset Statistics
Our IndustryASM dataset comprises of 4803 videos in total and the average video duration is 119 seconds. Therefore, the total dataset duration is 158 hours. These videos are distributed among 52 industry assembly processes or datasets, ranging from automotive manufacturing, electronic device manufacturing, medical device manufacturing to heating, ventilation, and air conditioning (HVAC) manufacturing. Each assembly process or dataset involves 12 steps or actions on average. We categorize each step into one of the following classes:
  • Moving: In this category, the worker(s) transfer parts or subassemblies within or between workstations, relocating items from one position to another within the workspace.
  • Assembling: Here, the worker(s) actively build or complete subassemblies by adding components or combining parts to form a complete unit.
  • Positioning: This involves the worker(s) placing parts onto a subassembly or mounting a subassembly in a designated position within the workstation.
  • Packaging: In this category, the worker(s) place subassemblies or components into boxes or other forms of packaging materials.
Examples of these 4 categories are presented in Fig. 3 of the main paper. We further classify each assembly process or dataset into one category which appears the most among all of its steps. For example, the Air_Cleaner dataset includes 10 steps (i.e., 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “moving” which appears the most among its 10 steps. We provide the statistics for each category as well as the entire IndustryASM dataset in Tab. 6.
(a) Complete Dataset
Number of videos 4803
Average video duration 119 seconds
Total dataset duration 158 hours
Average number of steps 12
Number of datasets 52
(b) Moving
Number of videos 383
Average video duration 176 seconds
Total dataset duration 19 hours
Average number of steps 11
Number of datasets 7
(c) Assembling
Number of videos 1671
Average video duration 144 seconds
Total dataset duration 67 hours
Average number of steps 14
Number of datasets 20
(d) Positioning
Number of videos 1696
Average video duration 109 seconds
Total dataset duration 51 hours
Average number of steps 13
Number of datasets 18
(e) Packaging
Number of videos 1053
Average video duration 74 seconds
Total dataset duration 22 hours
Average number of steps 10
Number of datasets 7
Table 6: Dataset statistics for IndustryASM.
A.3 Annotation Process
Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers. To ensure quality, two labelers are assigned for each video, i.e., one labeler provides the labels for the video, while another labeler checks the labels. If there are conflicts, both labelers discuss to fix them. In general, roughly 8% of the videos have conflicts and need fixing, resulting in an agreement rate of around 92% between labelers. We manually write the user instructions for the temporal action segmentation task and generate the ground truth responses by using the action names and action timestamps. Fig. 6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
Figure 6: Examples of our instruction and target response for the temporal action segmentation task on an IndustryASM video.
A.4 Implementation Details
We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab. 7. We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames. The frame encoder processes frames of sizes $224\times 224$ across $14\times 14$ patches. Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively. The output of the clip encoder contains a time-aware representation of each clip. We then concatenate the time-aware clip features according to their temporal order of appearance in the video. This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens. This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$. The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM. Projection layers are used to project the BiLSTM output into the LLM input space. The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
Hyperparameter Value
Number of clips $C$ 6
Number of frames sampled per clip $N^{c}_{f}$ 96
Frame encoder patch size 14$\times$14
Frame resolution 224$\times$224
Frame sampling type uniform
Number of epochs 7
Batch size 32
Learning rate 5e-5
Warmup learning rate 5e-6
Weight decay 0.01
Optimizer Adam
AdamW $\beta$ (0.9,0.999)
Scheduler LinearWarmupCosineLR
Q-Former window size 32
Stride 32
Nnumber of visual tokens per window 32
Number of layers in clip encoder 2
Number of layers in image encoder 12
Q-Former hidden size $N_{V}$ 768
BiLSTM input size 768
BiLSTM hidden size 768
LLM hidden size $N_{LLM}$ 4096
Table 7: Hyperparameter settings.
A.5 Comparison on Model Sizes
Tab. 8 compares the sizes of our TemporalVLM model, TimeChat [35], and LongVLM [49] in terms of the number of learnable parameters. Our TemporalVLM model includes a trainable BiLSTM module and projection layers, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat [35] and LongVLM [49] respectively. Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat [35]. In particular, TimeChat [35] is trained on the complete TimeIT [35] and Valley [27] datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download the YT-Temporal [54] dataset on time due to its large size and the restricted number of downloads by YouTube. of the TimeIT [35] and Valley [27] datasets.
Model Number of Trainable Parameters
TimeChat [35] 241 millions
LongVLM [49] 244 millions
TemporalVLM (Ours) 254 millions
Table 8: Comparison on model sizes.
A.6 General Video Understanding Results
Method Data CI DO CU TU C
Video-LLaMA [56] 10M 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT [28] 100k 2.50 2.57 2.69 2.16 2.20
Valley [27] 234k 2.43 2.13 2.86 2.04 2.45
BT-Adapter [24] 10M 2.68 2.69 3.27 2.34 2.46
LongVLM [49] 100k 2.76 2.86 3.34 2.39 3.11
TemporalVLM 100k 2.88 2.91 3.45 2.50 3.16
Table 9: General video understanding results. Best results are in bold, while second best ones are underlined.
Following LongVLM [49], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [28] in Tab. 9. The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C). It is evident from Tab. 9 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [49].
A.7 Qualitative Results
In addition to Figs. 4 and 5 in the main paper, we provide additional qualitative results in this section. In particular, Figs. 7, 8, and 9 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively. More importantly, qualitative results demonstrating the generalization abilities of TemporalVLM in the zero-shot setting are shown in Fig. 10. Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understaning tasks.
Figure 7: Qualitative examples highlighting the dense video captioning capabilities of TemporalVLM in the supervised setting. The model is asked to provide the timestamps of actions that occur in the videos along with brief descriptions of the actions.
Figure 8: Qualitative examples demonstrating the video highlight detection capabilities of TemporalVLM in the supervised setting. The model is given a video along with an action description. It is prompted to provide the frames that match the action description and their saliency scores.
Figure 9: Qualitative examples showing the temporal video grounding capabilities of TemporalVLM in the supervised setting. A video and a query is given to the model. It is prompted to provide the timestamps at which the query occurs. Magenta represents the ground truth and predicted timestamps for the first query, while Orange indicates those for the second query.
Figure 10: Examples of the generalization capabilities of TemporalVLM in the zero-shot setting. (Top) The model is prompted to provide the timestamps of actions and brief descriptions of actions for an egocentric video of furniture assembly. (Bottom) The model is provided with a documentary film. It is asked to predict the timestamps when the query happens. Blue denotes the ground truth and predicted timestamps.
A.8 Discussions
A.9 Limitations
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations. Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
A.10 Societal Impacts
Our time-sensitive video LLM could enable a variety of applications including optimizing and tracking industry assembly processes. More specifically, for assembly process optimization, industrial engineers could utilize TemporalVLM to automatically decompose a video recording of an assembly process into segments as well as generate a brief description for each segment. In addition, TemporalVLM could improve the performance of frontline workers by tracking their assembly process and notify them as soon as they miss a step. Nevertheless, we acknowledge that our time-aware video LLM could be misused for surveillance and monitoring of individuals, which emphasizes the importance of responsible AI principles to guide the use of this technology.

### v4

@@
 At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues.
-In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps into time-sensitive local features.
+In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps and fused across overlapping temporal windows into time-sensitive local features.
 Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation.
@@
 The above methods usually treat the entire video as a single clip [33, 13, 14, 36] and aggregate video tokens via pooling operation [14] and query aggregation [36], struggling to capture fine-grained information, especially for long videos.
-In this paper, we propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos.
+We propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos.
 Fig.
@@
 Our model includes two key architectural contributions for temporal reasoning in long videos.
-Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting time-aware local features from each clip.
+Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip.
 Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips.
@@
   • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos.
-Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into time-aware local features.
+Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into fused time-aware local features.
 The local features are then fed to a BiLSTM module to extract global features.
@@
 Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps.
-The time-aware clip encoder is then employed on each clip along with its timestamps for extracting time-aware local features, which capture local fine-grained cues within each clip.
+The time-aware clip encoder is employed on each clip along with its timestamps for extracting fused time-aware local features, which capture local fine-grained cues within each clip.
 Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips.
@@
 Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip.
-The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding time-aware local features.
-We adopt the time-aware frame encoder and sliding video Q-Former of TimeChat [35] as our time-aware clip encoder.
-The time-aware frame encoder uses a pre-trained image encoder [39] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features.
-These frame features are then fed to a sliding video Q-Former [35] for temporal modeling, yielding time-aware clip features for each clip.
-Recent works, e.g., LongVLM [49], leverage clip sampler and clip encoder for long video understanding but their encoders are not time-aware and hence are not effective for temporal reasoning.
+The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding fused time-aware local features.
+The time-aware frame encoder uses a pre-trained image encoder [40] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features $\mathbf{f}_{t}$.
+These frame features $\mathbf{f}_{t}$ are then fed to a video Q-Former  [36] across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 and overlap $o$ = 16 for temporal modeling.
+The window-wise outputs $\mathbf{V}_{i}$ are then concatenated into $\mathbf{S}$ :
+[…]
+$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}],$ (2)
+with $n$ denoting the number of overlapping windows within a short-term clip.
+We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$.
+Specifically, we apply multiheaded self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
+$\mathbf{C}^{(h)}$ $=\text{SoftMax}\left(\frac{\mathbf{Q}^{(h)}{\mathbf{K}^{(h)}}}{\sqrt{d}}\right)\mathbf{V}^{(h)},$ (3)
+$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}$ (4)
+We first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$.
+At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$.
+$\mathbf{C}$ contains the aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied.
+Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window context.
+Through Eq.
+4, each redundant time-aware token in $\mathbf{S}$ is enabled to compare and weigh its contextually dissimilar variant, aligning them into a fused temporal representation.
+This allows $\mathbf{C}$ to yield rich time-aware clip level information by leveraging diverse temporal views from across overlapping windows.
+Recent works, e.g., [50], leverage clip sampler and clip encoder for long video understanding but their encoders are not timeaware and hence are not effective for temporal reasoning.
 3.2 Bidirectional Long Short-Term Memory
@@
 We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video.
-We then pass the sequence of local features to a BiLSTM module for aggregating global features.
+We then pass the sequence of fused local features to a BiLSTM module for aggregating global features.
 Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
@@
 where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$.
-The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as:
-$\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}].$ (3)
+The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$.
 Please refer to the textbook [19] for a detailed description.
@@
 Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos.
-Unlike the recent LongVLM model [49], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec.
+Unlike LongVLM [50], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec.
 5.
@@
 Figure 3: Example videos with different camera viewpoints, actors, backgrounds, and activities from our IndustryASM dataset.
-We follow prior works [32, 13, 14, 35] to test TemporalVLM on dense video captioning, temporal video grounding, and video highlight detection.
-To further evaluate our model, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
-To our best knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers.
-Below we provide some details and annotation process.
+We introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
 4.1 Dataset Statistics
@@
 We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [57] checkpoints to initialize the video Q-Former.
+We employ pytorch’s [15] implementation of multiheaded attention in the fusion module of the clip encoder with 8 attention heads.
 The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer.
@@
 We use 6 clips and sample 96 frames from each clip.
-We fine-tune our model on a subset of TimeIT [35] and Valley [27] for 7 epochs on a single 8-A100 (80 GB) machine.
-The parameters of the LLM and image encoder remain frozen throughout the fine-tuning.
+We fine-tune our model on a subset of TimeIT [36] and Valley [28] on a single 8-A100 (80 GB) machine.
+The parameters of the LLM and image encoder remain frozen throughout fine-tuning.
 Please see our supplementary material for more details.
@@
 Best results are in bold, while second best ones are underlined.
-We first evaluate all methods in the zero-shot setting and present the results in Tab. 1.
-5.1.1 Dense Video Captioning Results
-For dense video captioning, it can be seen from Tab.
-1 that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score.
+Dense Video Captioning Results.
+Tab.
+1 shows that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score.
 For example, we outperform TimeChat [36] by +0.3 on CIDEr and +0.5 on F1 Score.
@@
 Moreover, our approach obtains superior results than LongVLM [50] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
-5.1.2 Video Highlight Detection Results
+Video Highlight Detection Results.
 The video highlight detection results in Tab.
@@
 Though LongVLM [50] extracts multi-level features, the lack of a time-aware encoder hinders their results.
-5.1.3 Temporal Video Grounding Results
+Temporal Video Grounding Results.
 It is evident from Tab.
@@
 The results validate the benefits of our time-aware and multi-level features for temporal video grounding.
-Our time-aware clip encoder extracts fine-grained cues for temporal reasoning within each clip, while our BiLSTM module captures long-range temporal relationships between multiple clips to detect the given event within the video effectively.
-The lack of a time-aware encoder reduces the performance of LongVLM [49], while TimeChat [35] suffers from a lack of fine-grained information.
+Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM module to capture long-range temporal relationships between multiple clips to detect the given event effectively.
+The lack of a time-aware encoder affects LongVLM [50], while TimeChat [36] suffers from a lack of fine-grained information.
 5.2 Supervised Results
-We now evaluate all methods in the supervised setting.
-5.2.1 Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results
+Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results.
+We present the results of all methods in the supervised setting in Tab.
+2.
+Our TemporalVLM model achieves the best results across all tasks and metrics.
+As compared to the zero-shot results in Tab.
+1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task.
+However, we acknowledge that task-specific models such as MMN [49] and QD-DETR [30] may outperform generalist models such as TimeChat [36] and our TemporalVLM model due to their task-specific design paradigms, e.g., using a mutually matching network to directly model the relationship between the query text and video content or specialized tokens and loss functions during fine-tuning.
 Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
@@
 Best results are in bold, while second best ones are underlined.
-We present the results of all methods in the supervised setting in Tab.
-2.
-It can be seen from Tab.
-2 that our TemporalVLM model achieves the best results across all tasks and metrics.
-As compared to the zero-shot results in Tab.
-1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task.
-However, we acknowledge that task-specific models such as MMN [48] and QD-DETR [29] may outperform generalist models such as TimeChat [35] and our TemporalVLM model due to their task-specific design paradigms such as using a mutually matching network to directly model the relationship between the query text and video content and introducing specialized tokens and loss functions during fine-tuning.
-5.2.2 Temporal Action Segmentation Results
+Temporal Action Segmentation Results.
+We further test all methods on our IndustryASM dataset for temporal action segmentation.
+Since the activities in IndustryASM significantly differ from those in TimeIT [36] and Valley [28], we fine-tune all models on IndustryASM before evaluation.
+Tab.
+3 presents the results.
+TemporalVLM performs the best across all metrics.
+For example, it outperforms TimeChat [36] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [50] by +8.9 and +5.3 on Edit and Acc respectively.
+The results confirm the benefits of our fused time-aware and multi-level features.
 Model F1@{10,25,50} Edit Acc
@@
 Best results are in bold, while second best ones are underlined.
-We further test all methods on our IndustryASM dataset for temporal action segmentation.
-Since the activities in IndustryASM significantly differ from those in TimeIT [35] and Valley [27], we fine-tune all models on IndustryASM before evaluation.
-Tab.
-3 presents the results.
-It is evident from Tab.
-3 that TemporalVLM performs the best across all metrics.
-For example, it outperforms TimeChat [35] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [49] by +8.9 and +5.3 on Edit and Acc respectively.
-The results confirm the benefits of our time-aware and multi-level features for temporal action segmentation.
 Multi-Clip Encoder vs.
@@
 5.3 Ablation Results
-# Clips SODA_c CIDEr F1
-2 2.6 9.2 18.5
-4 3.1 11.3 18.9
-6 3.4 13.2 20.0
-Table 4: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
-Model # Layers SODA_c CIDEr F1
-Average Pooling 1 1.7 6.0 16.1
-Linear Layer 1 2.4 9.5 17.8
-Transformer 2 2.3 6.8 15.6
-LSTM 1 2.8 9.4 18.2
-BiLSTM 2 3.4 13.2 20.0
-Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
-Below we study the effects of the number of short-term clips and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [58].
-5.3.1 Impacts of the Number of Short-Term Clips
-Tab.
-4 presents the effects of the number of short-term clips on our TemporalVLM model.
-Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6.
-It is evident from Tab.
-4 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
-5.3.2 Impacts of the BiLSTM Module
+Below we study the effects of our overlapping sliding video Q-Former and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [59].
+Impacts of Overlapping Sliding Video Q-Former.
+We conduct experiments in Tab.
+4 to demonstrate the importance of fusing encoded frames across overlapping temporal windows in our clip encoder.
+Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows or removing the fusion module leads to performance drops.
+The results validate the design of our overlapping sliding video Q-Former module.
+Model SODA CIDEr F1
+No Overlap + No Fusion 2.3 11.0 17.3
+No Overlap + Fusion 2.0 9.5 16.2
+Overlap + No Fusion 2.8 11.7 18.9
+Overlap + Fusion 3.4 13.2 20.0
+Table 4: Effects of overlapping sliding video Q-Former. Best results are in bold, second best are underlined.
+Impacts of BiLSTM.
 To study the effects of BiLSTM, we replace it with various alternatives and report the results in Tab.
@@
 Finally, BiLSTM outperforms transformer [44], which we will discuss below.
+Model # Layers SODA_c CIDEr F1
+Average Pooling 1 1.7 6.0 16.1
+Linear Layer 1 2.4 9.5 17.8
+Transformer 2 2.3 6.8 15.6
+LSTM 1 2.8 9.4 18.2
+BiLSTM 2 3.4 13.2 20.0
+Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
 (Bi)LSTM vs.
@@
 Black represents background frames.
-In this section, we discuss some qualitative results.
-5.4.1 Dense Video Captioning Results
+Dense Video Captioning Results.
 Fig.
@@
 LongVLM [50] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions.
-In contrast, TemporalVLM produces considerably more accurate timestamps and captions.
-Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long video understanding.
-5.4.2 Temporal Action Segmentation Results
+In contrast, TemporalVLM produces much more accurate timestamps and captions.
+Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long videos.
+Temporal Action Segmentation Results.
 Fig.
@@
 LongVLM [50] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video.
-In contrast, TemporalVLM predicts action segments that are significantly closer to the ground truth.
+In contrast, TemporalVLM predicts action segments significantly closer to the ground truth.
 6 Conclusion
 We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos.
-Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features.
+Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts fused time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features.
 The extracted features not only are time-sensitive but also contain both local and global cues.
-In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes.
-Our dataset is recorded on factory floors and includes action and timestamp labels provided by industrial engineers for time and motion studies and temporal action segmentation evaluation.
+In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for time and motion studies and temporal action segmentation evaluation.
 Lastly, we conduct extensive experiments on TimeIT [36] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks.
@@
 A.1 and A.4 respectively.
-Next, Sec.
-A.5 compares the sizes of our model, TimeChat [35], and LongVLM [49], while Sec.
-A.6 evaluates our model for general video understanding.
+Next, we provide a comparison with specialized models such as Vid2Seq [53] and QD-DETR [30] in Sec.
+A.5, followed by demonstrating the effects of number of training epochs in Sec.
+A.6.
+Sec.
+A.7 compares the sizes of our model, TimeChat [36], and LongVLM [50], while Sec.
+A.8 assesses the impact of number of clips.
+Sec.
+A.9 evaluates our model for general video understanding.
 We then present several qualitative results in Sec.
-A.7, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results.
+A.10, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results.
 Finally, we discuss the limitations and societal impacts of our work in Sec.
-A.8.
+A.11.
 A.1 IndustryASM Details
@@
 Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively.
-The output of the clip encoder contains a time-aware representation of each clip.
+Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q$ and overlap $o$.
+We then fuse the concatenated window-wise outputs of the Q-former via multiheaded self attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former.
+The output of the clip encoder contains a fused time-aware representation of each clip.
 We then concatenate the time-aware clip features according to their temporal order of appearance in the video.
@@
 Scheduler LinearWarmupCosineLR
-Q-Former window size 32
-Stride 32
-Nnumber of visual tokens per window 32
+Q-Former window size $q$ 32
+Overlap $o$ 16
+Number of visual tokens per window 32
 Number of layers in clip encoder 2
 Number of layers in image encoder 12
+No of attention heads in fusion module 8
 Q-Former hidden size $N_{V}$ 768
+Fusion module hidden size $D$ 768
 BiLSTM input size 768
@@
 Table 7: Hyperparameter settings.
-A.5 Comparison on Model Sizes
+A.5 Comparison With Specialized Models
+Tab.
+8 shows generalist models underperform specialized models such as Vid2Seq  [53] and QD-DETr [30] whose advantages come from task-specific training data and model designs.
+In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR  [30] leads to superior performance in highlight detection tasks.
+In contrast our TemporalVLM is trained on a simpler language modeling loss on lesser epochs.
+Generalist models such as ours exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
+Model DVC VHD TVG
+SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
+Vid2Seq 7.9 47.1 27.3
+QD-DETR 38.9 62.4
+QD-DETR w/ Audio 39.0 62.9
+MMN 50.5 29.7
+VDI 52.3 31.4
+TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
+Table 8: Comparisons with specialized models.
+A.6 Effects of Number of Training Epochs
+Tab.
+9 shows the performance of our models across different training epochs.Since our model uses a learnable fusion layer in the clip encoder and a BiLSTM module to aggregate clip level features into a global representation, we observe the best results by training for 7 epochs.
+However, Tab.
+9 reports our results with 3 epochs, which underperform those with 7 epochs but still outperform those of TimeChat and LongVLM.
+Model DVC VHD TVG
+SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
+LongVLM (3 epochs) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
+TimeChat (3 epochs) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
+TemporalVLM (3 epochs) 3.2 12.9 19.7 23.9 42.3 50.0 25.9
+TemporalVLM (7 epochs) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
+Table 9: Effects of number of training epochs.
+A.7 Comparison on Model Sizes
 Tab.
@@
 of the TimeIT [36] and Valley [28] datasets.
+A.8 Impacts of Number of Short-Term Clips
+Tab.
+11 presents the effects of the number of short-term clips on our TemporalVLM model.
+Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6.
+It is evident from Tab.
+11 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
 Model Number of Trainable Parameters
@@
 LongVLM [50] 244 millions
-TemporalVLM (Ours) 254 millions
+TemporalVLM (Ours) 255 millions
 Table 10: Comparison on model sizes.
-A.6 General Video Understanding Results
+# Clips SODA_c CIDEr F1
+2 2.6 9.2 18.5
+4 3.1 11.3 18.9
+6 3.4 13.2 20.0
+Table 11: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
+A.9 General Video Understanding Results
 Method Data CI DO CU TU C
@@
 Following LongVLM [50], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [29] in Tab.
-9.
+12.
 The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C).
@@
 12 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [50].
-A.7 Qualitative Results
+A.10 Qualitative Results
 In addition to Figs.
@@
 Blue denotes the ground truth and predicted timestamps.
-A.8 Discussions
-A.9 Limitations
+A.11 Discussions
+A.12 Limitations
 Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations.
 Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
-A.10 Societal Impacts
+A.13 Societal Impacts
 Our time-sensitive video LLM could enable a variety of applications including optimizing and tracking industry assembly processes.

### v5

@@
 Abstract
-This paper introduces TemporalVLM, a video large language model (video LLM) capable of effective temporal reasoning and fine-grained understanding in long videos.
-At the core, our approach includes a visual encoder for mapping a long-term input video into features which are time-aware and contain both local and global cues.
-In particular, it first divides the input video into short-term clips, which are jointly encoded with their timestamps and fused across overlapping temporal windows into time-sensitive local features.
+We introduce TemporalVLM, a video large language model (video LLM) for temporal reasoning and fine-grained understanding in long videos.
+Our approach includes a visual encoder for mapping a long-term video into features which are time-aware and contain both local and global cues.
+It first divides an input video into short-term clips, which are jointly encoded with timestamps and fused across overlapping temporal windows into time-sensitive local features.
 Next, the local features are passed through a bidirectional long short-term memory (BiLSTM) module for global feature aggregation.
-The extracted time-aware and multi-level features are important for accurate temporal reasoning and fine-grained understanding in long videos.
-Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, which consists of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation.
-Finally, extensive experiments on datasets of long videos, including TimeIT and IndustryASM, show that TemporalVLM achieves superior performance than previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
-To the best of our knowledge, our work is the first to incorporate LSTMs into video LLMs.
-Figure 1: Previous video large language models (a-c) usually are not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling operation (a, b) or query aggregation (c) for aggregating global semantic information.
-In contrast, our model (d) includes a time-aware clip encoder, which extracts time-aware fine-grained cues from short-term clips sampled from a long-term input video, and a bidirectional long short-term memory module, which captures long-range temporal dependencies across multiple clips.
-Our extracted time-aware and multi-level features are crucial for temporal reasoning in long videos.
+Moreover, to facilitate the evaluation of TemporalVLM, we present a large-scale long video dataset of industry assembly processes, namely IndustryASM, consisting of videos recorded on factory floors with actions and timestamps annotated by industrial engineers for time and motion studies and temporal action segmentation evaluation.
+Finally, extensive experiments show that TemporalVLM outperforms previous methods across temporal reasoning and fine-grained understanding tasks, i.e., dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
+To our best knowledge, our work is the first to incorporate LSTMs into video LLMs.
+Video LLMs for Temporal Reasoning in Long Videos
+Fawad J. Fateh    Umer Ahmed    Hamza Khan    M. Zeeshan Zia    Quoc-Huy Tran Retrocausal, Inc., Redmond, WA www.retrocausal.ai
 1 Introduction
@@
 {fawad,umer,hamza,zeeshan,huy}@retrocausal.ai.
+Figure 1: Video LLMs are often not time-sensitive (a, b), consider an input video as a single clip (a, c), and apply pooling (a, b) or query aggregation (c) for aggregating global features.
+Our model (d) includes a time-aware clip encoder for extracting time-aware fine-grained cues and a BiLSTM for capturing long-range temporal dependencies.
 Video temporal reasoning represents the process of reasoning about time and its passage in videos, with a focus on how events or actions happen and relate to each other in terms of time.
-Several video understanding applications require the ability of temporal reasoning.
-Examples include dense video captioning [46, 60, 53], which attempts to detect the timestamps and generate detailed descriptions of all the events or actions in an input video, and temporal action segmentation [23, 54, 2, 26], which aims to associate each frame of a long video capturing a complex activity with one of the action/sub-activity classes.
-Considerable efforts [53, 27, 30, 26] have been invested in developing separate models for solving individual tasks.
-These models usually have different architectures.
-Therefore, it is favorable to design a unified model for handling various tasks.
-The past few years have witnessed the impressive instruction following ability and remarkable comprehension and generation capability of large language models (LLMs) [1, 4, 41, 31, 42, 43], which have emerged as a universal agent for performing various tasks.
-Video LLMs [22, 24, 28, 29, 57, 16, 39] which incorporate video encoders with LLMs for video understanding tasks have been introduced.
-These methods often represent a video by a fixed number of visual tokens, leading to reduced performance with long videos, and encode video frames and their timestamps separately and hence struggle to perform temporal reasoning tasks.
-Recently, a few works [33, 13, 14, 36] have developed video LLMs with temporal reasoning abilities, e.g., TimeChat [36] proposes to vary the number of final video tokens based on the input video length and jointly encode video frames and their timestamps.
-The above methods usually treat the entire video as a single clip [33, 13, 14, 36] and aggregate video tokens via pooling operation [14] and query aggregation [36], struggling to capture fine-grained information, especially for long videos.
-We propose TemporalVLM, a video LLM capable of effective temporal reasoning and fine-grained understanding in long videos.
+Several video understanding applications require the ability of temporal reasoning, e.g., dense video captioning Wang et al.
+(2021); Zhu et al.
+(2022); Yang et al.
+(2023) and temporal action segmentation Li et al.
+(2020); Yi et al.
+(2021); Behrmann et al.
+(2022); Lu and Elhamifar (2024).
+Considerable efforts Yang et al.
+(2023); Luo et al.
+(2023a); Moon et al.
+(2023); Zhong et al.
+(2023); Lu and Elhamifar (2024); Lin et al.
+(2024) have been invested in developing models for solving individual tasks.
+These models often have different architectures.
+Thus, it is favorable to design a unified model for handling various tasks.
+The past few years have witnessed the impressive comprehension and generation capability of LLMs Achiam et al.
+(2023); Chiang et al.
+(2023); Taori et al.
+(2023); OpenAI (2024); Touvron et al.
+(2023a, b), which have emerged as a universal agent for performing various tasks.
+Video LLMs Li et al.
+(2023b); Liu et al.
+(2023); Luo et al.
+(2023b); Maaz et al.
+(2023); Zhang et al.
+(2023); Jin et al.
+(2024); Song et al.
+(2024) which incorporate video encoders with LLMs have been introduced.
+These methods often represent a video by a fixed number of tokens, yielding reduced performance with long videos, and encode frames and timestamps separately and hence struggle with temporal reasoning tasks.
+Please refer to the inferior results of these methods in Tab.
+1.
+Recently, a few works Qian et al.
+(2024); Huang et al.
+(2024a, b); Ren et al.
+(2024) have developed video LLMs with temporal reasoning abilities, e.g., TimeChat Ren et al.
+(2024) proposes to vary the number of tokens based on the video length and jointly encode frames and timestamps.
+The above methods usually treat the entire video as a single clip Qian et al.
+(2024); Huang et al.
+(2024a, b); Ren et al.
+(2024) and aggregate tokens via pooling operation Huang et al.
+(2024b) and query aggregation Ren et al.
+(2024), struggling to capture fine-grained details in long videos.
+We propose TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos.
 Fig.
-1 shows the differences in model architectures of TemporalVLM and previous works.
-Our model includes two key architectural contributions for temporal reasoning in long videos.
+1 shows the comparisons with prior works.
+Our model includes two architectural contributions.
 Firstly, we divide a long-term input video into multiple short-term clips and introduce a time-sensitive clip encoder for extracting fused time-aware local features from each clip.
-Secondly, we adopt a bidirectional long short-term memory (BiLSTM) module which takes all the local features as inputs and computes global features from multiple clips.
-As a consequence, our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos.
+Secondly, we adopt a BiLSTM module which takes all the local features as inputs and computes global features from multiple clips.
+Our features not only are time-sensitive but also contain both local fine-grained and global semantic information, which are crucial for temporal reasoning in long videos.
 Moreover, to further evaluate our model, we present IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation benchmarking and time and motion studies.
@@
 We convert the labels into chat samples with manually written instructions.
-To the best of our knowledge, it is the first large-scale video dataset recorded on factory floors and labeled by industrial engineers for temporal action segmentation evaluation and time and motion analyses.
-Lastly, extensive experiments on datasets of long videos, i.e., TimeIT [36] and IndustryASM, demonstrate that our TemporalVLM approach achieves superior results over previous methods across temporal reasoning and fine-grained understanding tasks, namely dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
+Lastly, extensive experiments demonstrate that TemporalVLM achieves superior results over previous methods.
 In summary, our contributions include:
-  • We develop TemporalVLM, a video LLM for temporal reasoning and fine-grained understanding in long videos.
-Given a long-term input video, we first sample short-term clips and encode each clip along with its timestamps into fused time-aware local features.
-The local features are then fed to a BiLSTM module to extract global features.
-Therefore, our features not only are time-sensitive but also capture effective local and global information.
-  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures.
-Our videos are recorded on factory floors with actions and timestamps labeled by industrial engineers for time and motion analyses and temporal action segmentation evaluation.
-We will make our dataset available publicly.
-  • Our evaluations show that TemporalVLM outperforms prior works across temporal reasoning and fine-grained understanding tasks.
-To our best knowledge, this is the first work to integrate LSTMs into video LLMs.
+  • We develop the first time-aware coarse-to-fine encoder, including a time-aware clip encoder (i.e., overlapping sliding video Q-Former) and a BiLSTM.
+By leveraging the modules, we tackle both fine-grained understanding and temporal reasoning in long videos.
+  • We present IndustryASM, a large-scale long video dataset of manufacturing assembly procedures. Our videos are recorded on factory floors and labeled by industrial engineers.
+  • TemporalVLM outperforms prior works on temporal reasoning and fine-grained understanding. To our best knowledge, this is the first work to blend LSTMs into video LLMs.
 2 Related Work
-Large Language Models.
-Over the past few years, LLMs [1, 4, 41, 31, 42, 43] have revolutionized the field of natural language processing.
-Initial methods employ encoder-decoder architectures [17, 35], whereas decoder-only models, e.g., GPT [34], demonstrate impressive performance and generalization in recent years.
-In addition, models with billions of parameters, e.g., GPT-4 [1], have been developed to push the boundary further, while instruction tuning techniques [32, 47, 48] have been widely used to enable models to generate contextually relevant responses.
-Here, we focus on video LLMs, which take videos as additional inputs and explore the use of LLMs for video understanding tasks.
 Video Large Language Models.
-Motivated by the impressive reasoning and generalization abilities of LLMs, video LLMs [22, 24, 28, 29, 57, 16, 39] have been developed for video understanding tasks.
-They typically include three key components: i) a pre-trained visual encoder (e.g., ViT [38]) to extract visual features, ii) a projection layer to map visual features into the text latent space of LLMs, and iii) a pre-trained LLM (e.g., LLaMA [42]) for generating responses.
-They mostly differ in their visual encoder, i.e., how visual features are extracted and aggregated from an input video.
-For example, VideoChat [22] extracts frame features via a visual transformer [38] and employs a query transformer (Q-Former) [21] to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA [57].
-The above methods usually map the video into a fixed number of visual tokens, yielding degrading performance with long videos, while encoding video frames and their timestamps separately and hence struggling with temporal reasoning tasks.
-To address these drawbacks, methods with temporal reasoning capabilities have been introduced, e.g., TimeChat [36] presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode video frames and their timestamps.
-These methods often consider the video as a single clip [28, 29, 22, 57] and aggregate video tokens via pooling operation [28, 29] and query aggregation [22, 57], overlooking fine-grained details.
-In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing local features from each clip.
-Also, we introduce a BiLSTM module for aggregating global features over multiple clips.
-To our best knowledge, our work is the first to exploit LSTMs for video LLMs.
+Video LLMs Li et al.
+(2023b); Liu et al.
+(2023); Luo et al.
+(2023b); Maaz et al.
+(2023); Zhang et al.
+(2023); Jin et al.
+(2024); Song et al.
+(2024) typically include a pre-trained visual encoder to extract visual features, a projection layer to map visual features into the text latent space of LLMs, and a pre-trained LLM for generating responses.
+They mostly differ in the visual encoder.
+VideoChat Li et al.
+(2023b) extracts frame features via a visual transformer Sharir et al.
+(2021) and employs a query transformer (Q-Former) Li et al.
+(2023a) to aggregate frame features into video features, while a video Q-Former is further included for temporal modeling in Video-LLaMA Zhang et al.
+(2023).
+The above methods usually map the video into a fixed number of tokens, yielding degrading performance with long videos, while encoding frames and timestamps separately and hence struggling with temporal reasoning.
+To address these drawbacks, methods with temporal reasoning capabilities are introduced, e.g., TimeChat Ren et al.
+(2024) presents a sliding video Q-Former to handle various video lengths and a time-aware frame encoder to jointly encode frames and timestamps.
+These methods often consider the video as a single clip Luo et al.
+(2023b); Maaz et al.
+(2023); Li et al.
+(2023b); Zhang et al.
+(2023) and aggregate tokens via pooling Luo et al.
+(2023b); Maaz et al.
+(2023) and query aggregation Li et al.
+(2023b); Zhang et al.
+(2023), overlooking fine-grained details.
+In this work, we divide the video into multiple clips and propose a time-aware clip encoder for capturing fused local features.
+Also, we integrate a BiLSTM module for aggregating global features.
 Video Temporal Reasoning.
-Temporal reasoning plays an important role in many video understanding tasks, e.g., dense video captioning [46, 60, 53], temporal video grounding [49, 27], video highlight detection [20, 30], and temporal action segmentation [23, 54, 2, 26].
-Previous works [53, 27, 30, 26] often focus on designing separate models for tackling individual tasks, while LLM-based models [33, 13, 14, 36] capable of handling multiple tasks have emerged recently.
+Temporal reasoning plays an important role in video understanding tasks, e.g., dense video captioning Wang et al.
+(2021); Zhu et al.
+(2022); Yang et al.
+(2023), temporal video grounding Wang et al.
+(2022); Luo et al.
+(2023a), video highlight detection Lei et al.
+(2021); Moon et al.
+(2023), and temporal action segmentation Li et al.
+(2020); Yi et al.
+(2021); Behrmann et al.
+(2022); Lu and Elhamifar (2024).
+Prior works Yang et al.
+(2023); Luo et al.
+(2023a); Moon et al.
+(2023); Lu and Elhamifar (2024) often focus on designing separate models for tackling individual tasks, while LLM-based models Qian et al.
+(2024); Huang et al.
+(2024a, b); Ren et al.
+(2024) capable of handling multiple tasks have emerged recently.
 Our TemporalVLM model belongs to the second group.
-Also, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion analyses.
 Long Video Understanding.
-There exist a few challenges in long video understanding such as complex spatial-temporal relationships and redundant information.
-Several methods have been developed for long video understanding, including efficient architectures [6, 18], temporal pooling/aggregation [37, 51], and dynamic clip selection [9, 10].
-For vision-language understanding tasks, methods based on temporal alignment [11, 3] have been introduced, while memory techniques [52, 58] are often used in dense video prediction tasks.
-Recently, Weng et al.
-[50] propose LongVLM for long video understanding.
-It divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features from multiple clips.
+Challenges in long video understanding include complex spatial-temporal relationships and redundant information.
+Long video understanding methods have been developed via efficient architectures Donahue et al.
+(2015); Kondratyuk et al.
+(2021), temporal pooling/aggregation Sener et al.
+(2020); Wu and Krahenbuhl (2021), and clip selection Ghodrati et al.
+(2021); Gowda et al.
+(2021).
+For vision-language understanding tasks, methods based on temporal alignment Han et al.
+(2022); Buch et al.
+(2022) and memory Wu et al.
+(2022); Zhao et al.
+(2023) have been introduced.
+Recently, LongVLM Weng et al.
+(2024) divides the video into clips and employs a merging module for extracting local features and pooling operation for computing global features.
 It does not utilize timestamps, which are crucial for temporal reasoning.
-In contrast, our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder, yielding time-aware features.
-Moreover, we employ a learnable BiLSTM module for aggregating global features from multiple clips.
-3 TemporalVLM: A Video LLM for Temporal Reasoning in Long Videos
-Figure 2: Given a long-term input video with timestamps, we divide it into short-term clips with timestamps.
-The time-aware clip encoder is employed on each clip along with its timestamps for extracting fused time-aware local features, which capture local fine-grained cues within each clip.
-Next, the bidirectional long short-term memory module is adopted for computing global features, which capture global temporal relationships across multiple clips.
-Lastly, the large language model takes video and text tokens as inputs and generates responses.
-Below we describe our first contribution, TemporalVLM, a video LLM for effective temporal reasoning in long videos.
-Two novel components are proposed in TemporalVLM: i) a time-aware clip encoder for extracting time-aware local features from a short-term clip, and ii) a BiLSTM module for computing global features from the local features.
-We provide an overview of our TemporalVLM model in Fig.
-2.
-3.1 Time-Aware Clip Encoder
-Video LLMs often consider a long-term input video as a single clip and uniformly sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat [36] and $N^{v}_{f}=100$ frames in VTimeLLM [13].
-Moreover, prior works usually use query aggregation, e.g., TimeChat [36], and pooling operation, e.g., Video-ChatGPT [29], to aggregate video tokens.
-Thus, they often perform well with short-term videos but miss out important fine-grained information for long-term videos, leading to degrading performance.
-We propose a time-aware clip encoder (i.e., blue block in Fig.
-2) for extracting local fine-grained cues within a clip.
-Given a long-term input video, we first decompose it into $C$ short-term clips and uniformly sample a fixed number of frames $N^{c}_{f}$ from each clip.
-Here, we set $C=6$ clips and $N^{c}_{f}=96$ frames per clip.
-The sampled frames for each clip along with their corresponding timestamps are then passed to our time-aware clip encoder for simultaneously encoding the frame contents and their timestamps, yielding fused time-aware local features.
-The time-aware frame encoder uses a pre-trained image encoder [40] to obtain frame features, which are jointly encoded with timestamps via an image Q-Former [5], yielding time-aware frame features $\mathbf{f}_{t}$.
-These frame features $\mathbf{f}_{t}$ are then fed to a video Q-Former  [36] across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 and overlap $o$ = 16 for temporal modeling.
-The window-wise outputs $\mathbf{V}_{i}$ are then concatenated into $\mathbf{S}$ :
-[…]
-$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}],$ (2)
-with $n$ denoting the number of overlapping windows within a short-term clip.
+Our TemporalVLM model explicitly utilizes timestamps via a time-aware clip encoder.
+Moreover, we employ a learnable BiLSTM module for aggregating global features.
+Procedural Activity Datasets.
+Existing datasets usually focus on cooking activities, e.g., Epic-Kitchens Damen et al.
+(2018), are curated from online sources and hence produced with multiple shots, e.g., COIN Tang et al.
+(2019), and are recorded with toy objects and lab environments, e.g., Assembly101 Sener et al.
+(2022).
+Our IndustryASM dataset focuses on manufacturing assembly processes, is recorded on factory floors, and is labeled by industrial engineers, thereby capturing procedural activities in realistic and practical environments that current datasets have not addressed.
+3 TemporalVLM
+Figure 2: TemporalVLM includes two novel components: a time-aware clip encoder for extracting time-aware fine-grained cues and a BiLSTM module for capturing long-range temporal dependencies.
+An overview of TemporalVLM is shown in Fig. 2.
+3.1 Short-Term Temporal Reasoning
+Video LLMs often consider an input video as a single clip and sample a fixed number of frames $N^{v}_{f}$ from the video, e.g., $N^{v}_{f}=96$ frames in TimeChat Ren et al.
+(2024).
+Moreover, prior works usually use query aggregation, e.g., TimeChat Ren et al.
+(2024), or pooling operation, e.g., Video-ChatGPT Maaz et al.
+(2023), to aggregate video tokens.
+Thus, they perform well with short videos but miss out fine-grained information for long videos.
+We introduce a time-aware clip encoder for extracting fused local fine-grained cues within a clip.
+We first divide the long-term input video into $C=6$ short-term clips and sample $N^{c}_{f}=96$ from each clip.
+Sampled frames of a clip along with timestamps are then passed to our time-aware clip encoder for jointly encoding frame contents and timestamps, yielding fused time-aware local features.
+Particularly, the time-aware frame encoder uses a pre-trained image encoder Sun et al.
+(2023) to obtain frame features, which are jointly encoded with timestamps via an image Q-Former Dai et al.
+(2023), yielding time-aware frame features $\mathbf{f}_{t}$.
+Overlapping Sliding Video Q-Former.
+Frame features $\mathbf{f}_{t}$ are fed to a video Q-Former  Ren et al.
+(2024) across overlapping windows $\mathbf{W}_{i}$ of size $q$ = 32 frames and overlap $o$ = 16 frames.
+Window-wise outputs $\mathbf{V}_{i}$ are concatenated into $\mathbf{S}$ as:
+$\mathbf{V}_{i}$ $=\text{Video Q-Former}(\mathbf{W}_{i}),$ (1)
+$\mathbf{S}$ $=[\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{W}],$ (2)
+with $W$ denoting the number of overlapping windows within a clip.
 We propose a fusion module to align diverse temporal cues from redundant boundary tokens in $\mathbf{S}$.
-Specifically, we apply multiheaded self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
+Specifically, we apply multi-headed self-attention on $\mathbf{S}$, yielding $\mathbf{C}$ which fuses local contexts across multiple windows into a single context-aware embedding as:
 $\mathbf{C}^{(h)}$ $=\text{SoftMax}\left(\frac{\mathbf{Q}^{(h)}{\mathbf{K}^{(h)}}}{\sqrt{d}}\right)\mathbf{V}^{(h)},$ (3)
-$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}$ (4)
-We first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$.
+$\mathbf{C}$ $=[\mathbf{C}^{(1)},\mathbf{C}^{(2)},\dots,\mathbf{C}^{(H)}]\mathbf{W}^{O}.$ (4)
+Here, we first project $\mathbf{S}$ into $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ for each attention head $h$.
 At each head $h$, $\mathbf{C}^{(h)}$ represents the dot product attention between queries $\mathbf{Q}^{(h)}$ and keys $\mathbf{K}^{(h)}$ scaled over dimension $d$ followed by softmax weighted aggregation over $\mathbf{V}^{(h)}$.
-$\mathbf{C}$ contains the aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied.
-Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window context.
-Through Eq.
-4, each redundant time-aware token in $\mathbf{S}$ is enabled to compare and weigh its contextually dissimilar variant, aligning them into a fused temporal representation.
-This allows $\mathbf{C}$ to yield rich time-aware clip level information by leveraging diverse temporal views from across overlapping windows.
-Recent works, e.g., [50], leverage clip sampler and clip encoder for long video understanding but their encoders are not timeaware and hence are not effective for temporal reasoning.
-3.2 Bidirectional Long Short-Term Memory
-Our time-sensitive clip encoder is applied on each short-term clip to obtain time-aware local features, which capture useful fine-grained cues for temporal reasoning within the short-term clip.
+$\mathbf{C}$ contains aggregated outputs from all heads with final projection $\mathbf{W}^{O}$ applied.
+Using overlapping windows leads to spatially redundant time-aware tokens in $\mathbf{S}$, each with different local window contexts.
+This allows $\mathbf{C}$ to yield rich time-aware clip-level information by leveraging diverse temporal views from across overlapping windows.
+TimeChat’s Ren et al.
+(2024) video Q-Former uses non-overlapping windows and does not perform fusion, yielding inferior results as shown in Tab.
+4.
+3.2 Long-Term Temporal Reasoning
+Our time-sensitive clip encoder is applied on each short-term clip to obtain fused time-aware local features, which capture useful fine-grained cues for temporal reasoning within the clip.
 However, they are not effective for temporal reasoning over the long-term video, which requires the ability to capture long-range temporal dependencies.
-We introduce a BiLSTM module (i.e., green block in Fig.
-2) for computing global features across multiple clips.
-We first concatenate time-aware local features extracted from clips in the temporal order which the clips appear in the video.
-We then pass the sequence of fused local features to a BiLSTM module for aggregating global features.
-Our BiLSTM module follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward LSTM) and another processes the sequence in the reverse order (backward LSTM) as:
+Bidirectional Long Short-Term Memory.
+We introduce a BiLSTM module for computing global features across multiple clips.
+We first concatenate time-aware local features extracted from clips in the temporal order the clips appear in the video.
+We then pass the sequence of local features to a BiLSTM for aggregating global features.
+Our BiLSTM follows a standard architecture, including two LSTM networks: one processes the sequence in the original order (forward) and another processes the sequence in the reverse order (backward) as:
 $\mathbf{h}^{f}_{t}=\text{LSTM}(\mathbf{h}^{f}_{t-1},\mathbf{c}_{t}),$ (5)
 $\mathbf{h}^{b}_{t}=\text{LSTM}(\mathbf{h}^{b}_{t+1},\mathbf{c}_{t}),$ (6)
-where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward LSTM and backward LSTM respectively and $\mathbf{c}_{t}$ is the input at time step $t$.
-The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward LSTM and backward LSTM as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$.
-Please refer to the textbook [19] for a detailed description.
+where $\mathbf{h}^{f}_{t}$ and $\mathbf{h}^{b}_{t}$ denote the hidden states at time step $t$ of the forward and backward LSTMs respectively and $\mathbf{c}_{t}$ is the input at time step $t$.
+The final output $\mathbf{h}_{t}$ at time step $t$ is obtained by concatenating the outputs of the forward and backward LSTMs as $\mathbf{h}_{t}=[\mathbf{h}^{f}_{t},\mathbf{h}^{b}_{t}]$.
+Please refer to the textbook Learning (2016) for a detailed description.
 Our BiLSTM module utilizes both past and future information and is capable of capturing long-range temporal relationships in both forward and backward directions.
-As empirically observed in Sec.
-5.3, BiLSTM outperforms various alternatives, including average pooling, linear layer, traditional LSTM, and transformer [44].
-Our final features output by BiLSTM not only are time-sensitive but also contain both local fine-grained and global semantic information, which are important for temporal reasoning in long videos.
-Unlike LongVLM [50], which applies pooling operation to obtain global features, our BiLSTM module has learnable parameters and achieves better results as shown in Sec.
-5.
-The video tokens output by our BiLSTM module have dimensions of $(C*N^{c}_{f},~{}2*N_{V})$, while the LLM requires an input of dimensions $(N^{c}_{f},~{}N_{LLM})$.
-Here, $N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM.
+As observed in Tab.
+5, BiLSTM outperforms various alternatives, including average pooling, linear layer, LSTM, and transformer Vaswani et al.
+(2017).
+Unlike LongVLM Weng et al.
+(2024), which applies pooling to obtain global features, our BiLSTM module has learnable parameters and achieves better results.
+Advanced recurring models, e.g., state space Gu et al.
+(2021) and Mamba Gu and Dao (2024), may further improve the performance, which remains our future work.
+The video tokens output by our BiLSTM module are of size $(C\times N^{c}_{f},~2\times N_{V})$, while the LLM requires an input size of $(N^{c}_{f},~N_{LLM})$.
+$N_{V}$ and $N_{LLM}$ are the dimensions of the video tokens and the LLM latent space respectively, and $2$ is to account for the outputs of both forward and backward passes of BiLSTM.
 Thus, we pass the video tokens output by BiLSTM through projection layers to match the input dimensions required by the LLM.
 3.3 Large Language Model
-The LLM (i.e., yellow block in Fig.
-2) takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users.
-Video LLMs typically employ a two-stage training scheme: i) leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and ii) utilizing instruction data for instruction following to fine-tune the pre-trained model.
-Following TimeChat [36], we use the checkpoint of the open-source LLaMA-2 7B model [43] and perform instruction tuning only.
-We optimize the below loss function:
+The LLM takes as input the video tokens $\mathbf{X}^{v}$, query tokens $\mathbf{X}^{q}$, and generates responses $\mathbf{X}^{r}$ to users.
+Video LLMs typically employ a two-stage training: leveraging large-scale image/video-text pairs for vision-language alignment to pre-train the model, and utilizing instruction data to fine-tune the pre-trained model.
+Following TimeChat Ren et al.
+(2024), we use the checkpoint of the LLaMA-2 7B model Touvron et al.
+(2023b) and perform instruction tuning only.
+We optimize the below loss:
 $\mathcal{L}=-\log P_{\mathbf{\psi}}(\mathbf{X}^{r}|\mathbf{X}^{v},\mathbf{X}^{q})$ (7)
@@
 Here, $\mathbf{\psi}$ is the learnable parameters of TemporalVLM, $L$ is the response length, $x_{i}$ is the current predicted token, and $\mathbf{X}^{r}_{<i}$ is the prior tokens appearing before $x_{i}$ in the response.
-Also, we employ LoRA [12] for efficient fine-tuning.
-4 IndustryASM: A Large-Scale Long Video Dataset of Industry Assembly Processes
-Figure 3: Example videos with different camera viewpoints, actors, backgrounds, and activities from our IndustryASM dataset.
-We introduce IndustryASM, a large-scale long video dataset of industry assembly processes for temporal action segmentation and time and motion studies.
+4 IndustryASM
+Figure 3: Example IndustryASM videos with different camera viewpoints, actors, backgrounds, and activities.
 4.1 Dataset Statistics
-Our IndustryASM dataset includes 4803 videos with each video lasting 119 seconds on average, yielding 158 hours as the total dataset duration.
-In addition, it consists of 52 various industry assembly procedures with each assembly procedure having 12 steps on average.
+IndustryASM includes 4803 videos with each lasting 119 seconds on average, yielding 158 hours as the total dataset duration.
+In addition, it consists of 52 various industry assembly procedures with each having 12 steps on average.
 Example videos are shown in Fig.
 3.
-Due to space limits, we provide additional details of IndustryASM in our supplementary material.
+Due to space limits, we provide additional details in our supplementary material.
 4.2 Annotation Process
-Our industry assembly videos are manually annotated by industrial engineers into relevant framewise action segments.
+Our videos are annotated by industrial engineers into framewise action segments.
 The action labels contain the same step naming conventions as used on factory floors during manufacturing processes.
-To ensure quality, each video is annotated by two labelers, i.e., one labeler obtains the labels for the video, while another labeler checks the labels.
+To ensure quality, each video is annotated by two labelers, i.e., one obtains the labels for the video, while another checks the labels.
 In cases of conflicts, both labelers discuss to address them.
-Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of roughly 92% between labelers.
+Overall, around 8% of the videos have conflicts and require fixing, yielding an agreement rate of 92% between labelers.
 Given the framewise annotations we first convert them to timestamps by using the frame rate of the video.
-We manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using timestamps and action labels.
-Example instructions and responses are shown in Fig.
-S1 of our supplementary material.
+We then manually write the instructions designed for the action segmentation task to ensure quality and generate the answers by using the timestamps and action labels.
+Examples are included in our supplementary material.
 5 Experiments
 Tasks and Datasets.
-Our TemporalVLM model is evaluated on a few temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
-For fair comparisons with TimeChat [36], we fine-tune our model on a subset [^1] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube.
-of the TimeIT [36] and Valley [28] datasets.
-The combined dataset contains a total of 142K long videos with diverse activities.
-We follow TimeChat [36] to use 6 instructions (generated by GPT-4 [1]) per task and convert the video annotations into chat samples.
-We evaluate on YouCook2 [59] for dense video captioning, Charades-STA [8] for temporal video grounding, QVHighlights [20] for video highlight detection, and our IndustryASM dataset for temporal action segmentation.
+We evaluate TemporalVLM on dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation.
+For fair comparisons with TimeChat Ren et al.
+(2024), we fine-tune our model on a subset [^1] We could not download YT-Temporal Zellers et al.
+(2022) due to its large size and the restricted number of downloads.
+of TimeIT Ren et al.
+(2024) and Valley Luo et al.
+(2023b).
+The combined dataset contains 142K long videos with diverse activities.
+We follow TimeChat to use 6 instructions (generated by GPT-4 Achiam et al.
+(2023)) per task and convert video annotations into chat samples.
+We evaluate on YouCook2 Zhou et al.
+(2018) for dense video captioning, Charades-STA Gao et al.
+(2017) for temporal video grounding, QVHighlights Lei et al.
+(2021) for video highlight detection, and IndustryASM for temporal action segmentation.
+Note that only Tab.
+3 uses our dataset, all the remaining evaluations use existing datasets.
+Following prior works, we report the results from a single run.
 Implementation Details.
-We follow TimeChat [36] to use LLaMA-2 7B [43] as the LLM, and ViT-G/14 from EVA-CLIP [40] as the image encoder.
-We take InstructBLIP’s checkpoints to initialize the image Q-Former, and Video-LLaMA’s [57] checkpoints to initialize the video Q-Former.
-We employ pytorch’s [15] implementation of multiheaded attention in the fusion module of the clip encoder with 8 attention heads.
-The BiLSTM module includes a forward LSTM with one hidden layer and a backward LSTM with one hidden layer.
-The parameters of the BiLSTM module and projection layers are initialized randomly.
-The LLM is fine-tuned using LoRA [12] with rank 32.
-We use 6 clips and sample 96 frames from each clip.
-We fine-tune our model on a subset of TimeIT [36] and Valley [28] on a single 8-A100 (80 GB) machine.
-The parameters of the LLM and image encoder remain frozen throughout fine-tuning.
+We follow TimeChat Ren et al.
+(2024) to use LLaMA-2 7B Touvron et al.
+(2023b) as the LLM, and ViT-G/14 from EVA-CLIP Sun et al.
+(2023) as the image encoder.
+We take InstructBLIP’s Dai et al.
+(2023) checkpoints to initialize the image Q-Former, and Video-LLaMA’s Zhang et al.
+(2023) checkpoints to initialize the video Q-Former.
+We employ pyTorch’s Imambi et al.
+(2021) multiheaded attention in the fusion module.
+The BiLSTM module includes forward and backward LSTMs with one hidden layer each.
+The BiLSTM and projection layers are initialized randomly, while the LLM and image encoder remain frozen throughout fine-tuning.
+The LLM is fine-tuned using LoRA Hu et al.
+(2023).
+We fine-tune our model on an 8-A100 (80 GB) machine.
 Please see our supplementary material for more details.
 Competing Methods.
-We compare our model against state-of-the-art video LLMs, including VideoChat-Embed [22], Video-LLaMA [57], Valley [28], TimeChat [36], and LongVLM [50].
-Since LongVLM [50] is not trained on TimeIT [36] and Valley [28] and it is a multi-model pipeline (TimeChat [36] and TemporalVLM are end-to-end models), we replace our time-aware clip encoder and BiLSTM modules with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM model which we use for evaluation.
-Nevertheless, we compare with the original LongVLM model [50] on general video understanding in our supplementary material.
-Evaluation Metrics. We utilize different sets of metrics for evaluating temporal reasoning tasks:
-  • Dense video captioning: The overall story-level evaluation is done through SODA_c [7].
-The event localization performance is measured using F1 Score.
-CIDEr [45] is used to measure the quality of the generated captions.
-  • Temporal video grounding: To gauge the effectiveness of temporal video grounding, we use R@1(IoU = $\alpha$), which is the percentage of predictions with an intersection over union greater than $\alpha$ compared to the ground truth.
-  • Video highlight detection: We use mean average precision (mAP) and HIT@1 at IoU= 0.75 to evaluate the accuracy of the saliency scores for key moments in videos.
-  • Temporal action segmentation: We report the framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance.
-We use the intersection over union ratio to find the overlapping thresholds.
+We compare our model against VideoChat-Embed Li et al.
+(2023b), Video-LLaMA Zhang et al.
+(2023), Valley Luo et al.
+(2023b), TimeChat Ren et al.
+(2024), and LongVLM Weng et al.
+(2024).
+Since LongVLM is not trained on TimeIT Ren et al.
+(2024) and Valley Luo et al.
+(2023b) and it is a multi-model pipeline (TimeChat and TemporalVLM are end-to-end), we replace our time-aware clip encoder and BiLSTM with their local and global feature extraction modules respectively while keeping the same image encoder and LLM, yielding our reimplemented LongVLM for evaluation.
+Nevertheless, we compare with the original LongVLM on general video understanding in our supplementary material.
+Evaluation Metrics. We use the below metrics:
+  • Dense video captioning: We use SODA_c Fujita et al. (2020) for story evaluation, F1 Score for event localization, and CIDEr Vedantam et al. (2015) for measuring caption quality.
+[…]
+  • Video highlight detection: We use mean average precision (mAP) and HIT@1 (IoU= 0.75) to evaluate saliency scores of key moments.
+  • Temporal action segmentation: We use framewise accuracy, F1 scores at overlapping thresholds $\{0.1,0.25,0.5\}$, and Edit distance.
 5.1 Zero-Shot Results
-Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
-YouCook2 [59] QVHighlights [20] Charades-STA [8]
+Model Dense Captioning Highlight Detection Temporal Grounding
+YouCook2 QVHighlights Charades-STA
 SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
-Valley [28] 0.1 0.0 1.5 10.9 15.2 4.7 1.6
-Video-LLaMA [57] 0.0 0.0 0.1 11.3 15.6 2.7 1.2
-VideoChat-Embed [22] 0.2 0.6 3.4 13.1 18.1 3.2 1.4
-LongVLM [50] 0.8 2.5 12.3 11.1 15.0 13.9 6.1
-TimeChat [36] 1.2 3.4 12.6 14.5 23.9 27.9 12.3
-TemporalVLM (Ours) 1.2 3.7 13.1 16.4 31.3 30.1 13.2
+Valley Luo et al. (2023b) 0.1 0.0 1.5 10.9 15.2 4.7 1.6
+Video-LLaMA Zhang et al. (2023) 0.0 0.0 0.1 11.3 15.6 2.7 1.2
+VideoChat-Embed Li et al. (2023b) 0.2 0.6 3.4 13.1 18.1 3.2 1.4
+LongVLM Weng et al. (2024) 0.8 2.5 12.3 11.1 15.0 13.9 6.1
+TimeChat  Ren et al. (2024) 1.2 3.4 12.6 14.5 23.9 27.9 12.3
+TemporalVLM 1.2 3.7 13.1 16.4 31.3 30.1 13.2
 Table 1: Zero-short results.
-We evaluate all methods on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding.
-Best results are in bold, while second best ones are underlined.
+We test on YouCook2 for dense video captioning, QVHighlights for video highlight detection, Charades-STA for temporal video grounding.
+Best results are in bold, second best ones are underlined.
 Dense Video Captioning Results.
 Tab.
-1 shows that our TemporalVLM model achieves the best results across all three metrics, i.e., SODA_c, CIDEr, and F1 Score.
-For example, we outperform TimeChat [36] by +0.3 on CIDEr and +0.5 on F1 Score.
-This is likely because TimeChat [36] treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information.
+1 shows that TemporalVLM achieves the best results across all metrics.
+For example, we outperform TimeChat Ren et al.
+(2024) by +0.3 on CIDEr and +0.5 on F1 Score.
+This is likely because TimeChat treats the entire video as a single clip and aggregates video tokens via query aggregation, overlooking fine-grained information.
 In contrast, our extracted features, which capture both local fine-grained cues from each clip and global temporal dependencies across multiple clips, are effective for caption generation and event localization.
-Moreover, our approach obtains superior results than LongVLM [50] across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
+Our approach also obtains superior results than LongVLM Weng et al.
+(2024) across SODA_c (+0.4), CIDEr (+1.2), and F1 Score (+0.8), which confirms the importance of our time-aware clip encoder and BiLSTM for dense video captioning.
 Video Highlight Detection Results.
-The video highlight detection results in Tab.
-1 show that our TemporalVLM model obtains the best performance on both mAP and HIT@1.
-In particular, it surpasses TimeChat [36] by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM [50] by +5.3 at mAP and by +16.3 at HIT@1.
+From Tab.
+1, TemporalVLM obtains the best performance on both mAP and HIT@1.
+It surpasses TimeChat Ren et al.
+(2024) by +1.9 in mAP and +7.4 in HIT@1, while outperforming LongVLM Weng et al.
+(2024) by +5.3 at mAP and by +16.3 at HIT@1.
 These results suggest the effectiveness of our model for video highlight detection.
-By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their ranks of importance within the video against a given query.
-Though LongVLM [50] extracts multi-level features, the lack of a time-aware encoder hinders their results.
+By aggregating global features from time-aware local features, it identifies the most relevant timestamps and assigns them saliency scores based on their relative importance within the video for a given query.
+Though LongVLM extracts multi-level features, the lack of a time-aware encoder hinders its results.
 Temporal Video Grounding Results.
 It is evident from Tab.
-1 that TemporalVLM achieves the best performance for temporal video grounding across R@1 (IoU=0.5) and R@1 (IoU=0.7).
-Specifically, it outperforms TimeChat [36] by +2.2 and +0.9 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively, and LongVLM [50] by +16.2 and +7.1 on R@1 (IoU=0.5) and R@1 (IoU=0.7) respectively.
+1 that TemporalVLM achieves the best performance across R@1(IoU=0.5) and R@1(IoU=0.7).
+It outperforms TimeChat Ren et al.
+(2024) by +2.2 and +0.9 on R@1(IoU=0.5) and R@1(IoU=0.7) respectively, and LongVLM Weng et al.
+(2024) by +16.2 and +7.1 on R@1(IoU=0.5) and R@1(IoU=0.7) respectively.
 The results validate the benefits of our time-aware and multi-level features for temporal video grounding.
-Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM module to capture long-range temporal relationships between multiple clips to detect the given event effectively.
-The lack of a time-aware encoder affects LongVLM [50], while TimeChat [36] suffers from a lack of fine-grained information.
+Our clip encoder extracts fine-grained cues for temporal reasoning within each clip via fusion, which enables our BiLSTM to capture long-range temporal relationships across clips to detect the given event effectively.
+The lack of a time-aware encoder affects LongVLM, while TimeChat suffers from a lack of fine-grained cues.
 5.2 Supervised Results
 Dense Video Captioning, Video Highlight Detection, and Temporal Video Grounding Results.
-We present the results of all methods in the supervised setting in Tab.
-2.
-Our TemporalVLM model achieves the best results across all tasks and metrics.
+Tab.
+2 presents the results in the supervised setting.
+TemporalVLM performs the best across all tasks and metrics.
 As compared to the zero-shot results in Tab.
-1, we see a further increase in performance, e.g., +24.3 on R@1 (IoU=0.5) in the temporal video grounding task and +9.5 on CIDEr in the dense video captioning task.
-However, we acknowledge that task-specific models such as MMN [49] and QD-DETR [30] may outperform generalist models such as TimeChat [36] and our TemporalVLM model due to their task-specific design paradigms, e.g., using a mutually matching network to directly model the relationship between the query text and video content or specialized tokens and loss functions during fine-tuning.
-Model Dense Video Captioning Video Highlight Detection Temporal Video Grounding
-YouCook2 [59] QVHighlights [20] Charades-STA [8]
+1, we see an increase in performance, e.g., +24.3 on R@1(IoU=0.5) in temporal video grounding and +9.5 on CIDEr in dense video captioning.
+However, we acknowledge that task-specific models, e.g., MMN Wang et al.
+(2022) and QD-DETR Moon et al.
+(2023), may outperform generalist models, e.g., TimeChat Ren et al.
+(2024) and TemporalVLM, due to their task-specific designs.
+Due to space limits, we provide a comparison with task-specific models in our supplementary material.
+Model Dense Captioning Highlight Detection Temporal Grounding
+YouCook2 QVHighlights Charades-STA
 SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
-LongVLM [50] 2.3 8.1 16.9 16.0 22.5 27.2 11.9
-TimeChat [36] 3.1 10.3 19.5 21.7 37.9 46.7 23.7
-TemporalVLM (Ours) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
+LongVLM Weng et al. (2024) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
+TimeChat Ren et al. (2024) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
+TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
 Table 2: Supervised results.
-We first fine-tune all models on TimeIT [36] and Valley [28], and then perform task-specific fine-tuning and evaluation on YouCook2 [59] for dense video captioning, QVHighlights [20] for video highlight detection, and Charades-STA [8] for temporal video grounding.
+We first fine-tune all models on TimeIT and Valley, and then perform task-specific fine-tuning and evaluation on YouCook2 for dense video captioning, QVHighlights for video highlight detection, and Charades-STA for temporal video grounding.
 Best results are in bold, while second best ones are underlined.
 Temporal Action Segmentation Results.
-We further test all methods on our IndustryASM dataset for temporal action segmentation.
-Since the activities in IndustryASM significantly differ from those in TimeIT [36] and Valley [28], we fine-tune all models on IndustryASM before evaluation.
+We now evaluate on our IndustryASM dataset.
+Since the activities in IndustryASM significantly differ from those in TimeIT Ren et al.
+(2024) and Valley Luo et al.
+(2023b), we fine-tune all models on IndustryASM before evaluation.
 Tab.
@@
 TemporalVLM performs the best across all metrics.
-For example, it outperforms TimeChat [36] by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM [50] by +8.9 and +5.3 on Edit and Acc respectively.
-The results confirm the benefits of our fused time-aware and multi-level features.
+For example, it outperforms TimeChat Ren et al.
+(2024) by +23.5 and +6.9 on Edit and Acc respectively, and LongVLM Weng et al.
+(2024) by +8.9 and +5.3 on Edit and Acc respectively.
+The results confirm the benefits of our time-aware and multi-level features.
 Model F1@{10,25,50} Edit Acc
-TimeChat [36] {11.1, 8.5, 4.4} 32.7 46.0
-LongVLM [50] {17.1, 13.3, 7.3} 47.3 47.6
-TemporalVLM (Ours) {22.3, 18.3, 11.1} 56.2 52.9
-Table 3: Temporal action segmentation results in the supervised setting on IndustryASM.
+LongVLM Weng et al. (2024) {17.1, 13.3, 7.3} 47.3 47.6
+TimeChat Ren et al. (2024) {11.1, 8.5, 4.4} 32.7 46.0
+TemporalVLM {22.3, 18.3, 11.1} 56.2 52.9
+Table 3: Temporal action segmentation results in supervised setting on IndustryASM.
 All models are fine-tuned on IndustryASM before evaluation.
@@
 Time-Aware Encoder.
-Multi-clip encoders in LongVLM [50] and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat [36] on temporal action segmentation in Tab.
+Multi-clip encoders in LongVLM Weng et al.
+(2024) and TemporalVLM are crucial to capturing fine-grained cues, leading to superior results over TimeChat Ren et al.
+(2024) on temporal action segmentation in Tab.
 3.
-In contrast, time-aware encoders in TimeChat [36] and TemporalVLM are important to event localization and temporal modeling, resulting in better performance than LongVLM [50] on dense video captioning, video highlight detection, and temporal video grounding in Tabs.
+In contrast, time-aware encoders in TimeChat and TemporalVLM are important to event localization and temporal modeling, yielding better results than LongVLM on dense video captioning, video highlight detection, and temporal video grounding in Tabs.
 1 and 2.
-Our TemporalVLM encoder is both multi-clip and time-aware, yielding the best performance across all tasks.
+TemporalVLM is both multi-clip and time-aware.
 5.3 Ablation Results
-Below we study the effects of our overlapping sliding video Q-Former and BiLSTM. For this purpose, we fine-tune multiple configurations of TemporalVLM on YouCook2 [59].
+We fine-tune several variants of TemporalVLM on YouCook2 Zhou et al. (2018) to study our overlapping sliding video Q-Former and BiLSTM.
 Impacts of Overlapping Sliding Video Q-Former.
-We conduct experiments in Tab.
-4 to demonstrate the importance of fusing encoded frames across overlapping temporal windows in our clip encoder.
-Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows or removing the fusion module leads to performance drops.
-The results validate the design of our overlapping sliding video Q-Former module.
+Tab.
+4 demonstrates the importance of fusing encoded frames across overlapping temporal windows in our clip encoder.
+Our model with fusion across overlapping windows achieves the best results, whereas using non-overlapping windows and removing the fusion module as in TimeChat’s Ren et al.
+(2024) design lead to performance drops.
+The results validate our design.
 Model SODA CIDEr F1
@@
 From Tab.
-5, BiLSTM outperforms average pooling, which is not trained and employed in prior works [29, 50], and traditional LSTM, which only performs the forward pass and hence relies only on past cues.
+5, BiLSTM outperforms average pooling, which is not trained and employed in prior works Maaz et al.
+(2023); Weng et al.
+(2024), and traditional LSTM, which only performs the forward pass and hence relies only on past cues.
 In contrast, BiLSTM is learnable and conducts both forward and backward passes to utilize both past and future cues.
-Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers [57, 36] alone, leading to worse results.
-This highlights the importance of BiLSTM even when video Q-Formers [57, 36] are being used.
-Finally, BiLSTM outperforms transformer [44], which we will discuss below.
+Moreover, we replace BiLSTM with linear layer to examine the impacts of video Q-Formers Zhang et al.
+(2023); Ren et al.
+(2024) alone, leading to worse results.
+This highlights the importance of BiLSTM even when video Q-Formers are being used.
+Finally, BiLSTM outperforms transformer Vaswani et al.
+(2017), which we will discuss below.
 Model # Layers SODA_c CIDEr F1
@@
 LSTM 1 2.8 9.4 18.2
+LSTM 2 2.9 9.3 18.5
 BiLSTM 2 3.4 13.2 20.0
-Table 5: Effects of the BiLSTM module. Best results are in bold, while second best ones are underlined.
+Table 5: Effects of BiLSTM module. Best results are in bold, while second best ones are underlined.
 (Bi)LSTM vs.
 Transformer.
-The permutation-invariant nature of the self-attention mechanism in transformer [44] leads to loss in temporal information of the input sequence.
-Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs [56].
-In contrast, (Bi)LSTM processes data one by one and preserves the order of the sequence, making them better suited for temporal modeling.
-Similar to [56], we also observe in Tab.
-5 that linear layer outperforms transformer [44].
+The permutation-invariant nature of the self-attention mechanism in transformer Vaswani et al.
+(2017) leads to loss in temporal information of the input sequence.
+Though using positional encoding alleviates this issue, it is inevitable that some temporal information loss occurs Zeng et al.
+(2023).
+In contrast, (Bi)LSTM processes data one by one and preserves the sequence order, making them better suited for temporal modeling.
+Similar to Zeng et al.
+(2023), we also observe in Tab.
+5 that linear layer outperforms transformer Vaswani et al.
+(2017).
 5.4 Qualitative Results
-Figure 4: Example dense video captioning results in the zero-shot setting on a YouCook2 [59] video. Inaccuracies are shown in red.
-Figure 5: Example temporal action segmentation results in the supervised setting on an IndustryASM video.
-Colored segments represent predicted actions with a particular color denoting the same action across all models.
-Black represents background frames.
+Figure 4: Dense video captioning in zero-shot setting on YouCook2. Red denotes inaccuracies.
+Figure 5: Temporal action segmentation in supervised setting on IndustryASM. Black denotes background.
 Dense Video Captioning Results.
 Fig.
-4 shows example results of dense video captioning in the zero-shot setting on a YouCook2 [59] video.
-TimeChat [36] fails to predict any correct timestamps and tends to hallucinate.
-LongVLM [50] predicts some timestamps correctly and generates more accurate captions, although it still hallucinates objects not present in the video and produces inaccurate/repeated captions.
+4 shows example results in the zero-shot setting on a YouCook2 Zhou et al.
+(2018) video.
+TimeChat Ren et al.
+(2024) fails to predict any correct timestamps and tends to hallucinate.
+LongVLM Weng et al.
+(2024) yields better results but still hallucinates objects not present in the video and produces inaccurate captions.
 In contrast, TemporalVLM produces much more accurate timestamps and captions.
-Also, it is the only model that provides captions till the end of the video, showing the effectiveness of the model for long videos.
+Also, it is the only model that provides captions till the end of the video, showing the effectiveness for long video understanding.
 Temporal Action Segmentation Results.
 Fig.
-5 illustrates example results of temporal action segmentation in the supervised setting on an IndustryASM video.
-TimeChat [36] falls short in segmenting the video correctly and hallucinates actions not present in the video.
-LongVLM [50] fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video.
-In contrast, TemporalVLM predicts action segments significantly closer to the ground truth.
+5 presents example results in the supervised setting on an IndustryASM video.
+TimeChat Ren et al.
+(2024) fails to segment the video and hallucinates actions not present in the video.
+LongVLM Weng et al.
+(2024) fares better in terms of the actions detected in the video but struggles with predicting the action segments and hallucinations further into the video.
+In contrast, TemporalVLM predicts the action segments notably closer to the ground truth.
+Supplementary Material. Due to space constraints, we include additional details, evaluations, and discussions in our supplementary material.
 6 Conclusion
-We propose TemporalVLM, a video LLM for effective temporal reasoning and fine-grained understanding in long videos.
-Our approach consists of two main architectural contributions: i) a time-aware clip encoder which extracts fused time-aware local features from multiple short-term clips sampled from a long-term input video and ii) a BiLSTM module which computes global features from the local features.
-The extracted features not only are time-sensitive but also contain both local and global cues.
-In addition, we introduce IndustryASM, a large-scale long video dataset of industry assembly processes for time and motion studies and temporal action segmentation evaluation.
-Lastly, we conduct extensive experiments on TimeIT [36] and IndustryASM to demonstrate our superior results over previous methods across temporal reasoning and fine-grained understanding tasks.
-To our best knowledge, this is the first work to incorporate LSTMs into video LLMs.
-Appendix A Supplementary Material
-In this supplementary material, we first provide additional details of our IndustryASM dataset and our TemporalVLM implementation in Secs.
-A.1 and A.4 respectively.
-Next, we provide a comparison with specialized models such as Vid2Seq [53] and QD-DETR [30] in Sec.
-A.5, followed by demonstrating the effects of number of training epochs in Sec.
-A.6.
-Sec.
-A.7 compares the sizes of our model, TimeChat [36], and LongVLM [50], while Sec.
-A.8 assesses the impact of number of clips.
-Sec.
-A.9 evaluates our model for general video understanding.
+We propose TemporalVLM for temporal reasoning and fine-grained understanding in long videos.
+Our approach includes a time-aware clip encoder for extracting fused time-aware local features and a BiLSTM module for global feature aggregation.
+Extracted features not only are time-sensitive but also contain both local and global cues.
+Moreover, we present IndustryASM, a long video dataset of industry assembly processes.
+Lastly, extensive experiments demonstrate our superior results over prior works across temporal reasoning and fine-grained understanding tasks.
+To our best knowledge, this is the first work to blend LSTMs into video LLMs.
+Our future work will explore advanced recurring models Gu et al.
+(2021); Gu and Dao (2024).
+Supplementary Material
+In this supplementary material, we first provide the additional details of our IndustryASM dataset (including dataset statistics, annotation process, maintenance and release plans) and our TemporalVLM implementation in Secs.
+A and B respectively.
+Next, we provide the comparisons with specialized models (e.g., Vid2Seq Yang et al.
+(2023) and QD-DETR Moon et al.
+(2023)) in Sec.
+C, followed by demonstrating the effects of number of training epochs in Sec.
+D.
+We then evaluate our model for general video understanding in Sec.
+E.
+Furthermore, Sec.
+F compares the sizes of our model against TimeChat Ren et al.
+(2024) and LongVLM Weng et al.
+(2024), while we report the run time comparisons in Sec.
+G to evaluate the efficiency of our model.
+Moreover, Sec.
+H assesses the impacts of number of clips, while we study the effects of employing a more advanced LLM in our model in Sec.
+I.
 We then present several qualitative results in Sec.
-A.10, including dense video captioning results, video highlight detection results, temporal video grounding results, and more importantly, generalization results.
-Finally, we discuss the limitations and societal impacts of our work in Sec.
-A.11.
-A.1 IndustryASM Details
-A.2 Dataset Statistics
+J, including dense video captioning, video highlight detection, temporal video grounding, and more importantly, generalization results.
+Finally, we discuss the limitations and societal impacts of our work in Secs.
+K and L respectively.
+Appendix A IndustryASM Details
+A.1 Dataset Statistics
 Our IndustryASM dataset comprises of 4803 videos in total and the average video duration is 119 seconds.
@@
 These videos are distributed among 52 industry assembly processes or datasets, ranging from automotive manufacturing, electronic device manufacturing, medical device manufacturing to heating, ventilation, and air conditioning (HVAC) manufacturing.
+Participants in our dataset are salaried factory workers based in the US.
+Informed consent is obtained from all participants in accordance with standard operating procedures prior to data collection.
+As an additional privacy safeguard, all faces and other personally identifiable information (PII) are blurred or otherwise anonymized in the recorded data.
 Each assembly process or dataset involves 12 steps or actions on average.
-We categorize each step into one of the following classes:
+We first categorize each step into one of the following classes:
   • Moving: In this category, the worker(s) transfer parts or subassemblies within or between workstations, relocating items from one position to another within the workspace.
-  • Assembling: Here, the worker(s) actively build or complete subassemblies by adding components or combining parts to form a complete unit.
+  • Assembling: The worker(s) build or complete subassemblies by adding components or combining parts to form a complete unit.
   • Positioning: This involves the worker(s) placing parts onto a subassembly or mounting a subassembly in a designated position within the workstation.
   • Packaging: In this category, the worker(s) place subassemblies or components into boxes or other forms of packaging materials.
-Examples of these 4 categories are presented in Fig.
+Next, we further classify each assembly process or dataset into one of the above categories which appears the most among all of its steps.
+For example, the Air_Cleaner dataset includes 10 steps (namely, 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “Moving” which appears the most among its 10 steps.
+We provide the statistics for each category of the IndustryASM dataset in Fig.
+6.
+Finally, examples of these 4 categories are presented in Fig.
 3 of the main paper.
-We further classify each assembly process or dataset into one category which appears the most among all of its steps.
-For example, the Air_Cleaner dataset includes 10 steps (i.e., 4 moving steps, 3 assembling steps, and 3 positioning steps), and hence it is classified as “moving” which appears the most among its 10 steps.
-We provide the statistics for each category as well as the entire IndustryASM dataset in Tab.
-6.
-(a) Complete Dataset
-Number of videos 4803
-Average video duration 119 seconds
-Total dataset duration 158 hours
-Average number of steps 12
-Number of datasets 52
-(b) Moving
-Number of videos 383
-Average video duration 176 seconds
-Total dataset duration 19 hours
-Average number of steps 11
-Number of datasets 7
-(c) Assembling
-Number of videos 1671
-Average video duration 144 seconds
-Total dataset duration 67 hours
-Average number of steps 14
-Number of datasets 20
-(d) Positioning
-Number of videos 1696
-Average video duration 109 seconds
-Total dataset duration 51 hours
-Average number of steps 13
-Number of datasets 18
-(e) Packaging
-Number of videos 1053
-Average video duration 74 seconds
-Total dataset duration 22 hours
-Average number of steps 10
-Number of datasets 7
-Table 6: Dataset statistics for IndustryASM.
-A.3 Annotation Process
+In addition, Fig.
+7 presents more examples to better illustrate the diverse camera viewpoints, actors, backgrounds, and activities in our IndustryASM dataset that are common in manufacturing settings.
+Figure 6: Dataset statistics for our IndustryASM dataset.
+Figure 7: Example videos with different camera viewpoints, actors, backgrounds, and activities from our IndustryASM dataset.
+Figure 8: Examples of our instruction and target response for the temporal action segmentation task on an IndustryASM video.
+A.2 Annotation Process
 Our IndustryASM videos are annotated with action names and action timestamps by industrial engineers.
@@
 Fig.
-6 illustrates examples of our instruction and ground truth response for the temporal action segmentation on an IndustryASM video.
-Figure 6: Examples of our instruction and target response for the temporal action segmentation task on an IndustryASM video.
-A.4 Implementation Details
-We present a detailed list of hyperparameter values used in tuning our TemporalVLM model in Tab.
-7.
+8 illustrates examples of our instruction and ground truth response for temporal action segmentation on an IndustryASM video.
+A.3 Maintenance and Release Plans
+We will host our IndustryASM dataset on Hugging Face and assign a permanent DOI to it to ensure long-term accessibility.
+We will release version updates to incorporate any corrections and add additional data/annotations if and when available.
+These updates will follow a versioning scheme (e.g., v1.0) and there will be documentations on any changes in the release notes.
+We will track and respond to user reported issues and errors on our GitHub repository.
+We will release the entire dataset upon acceptance of this paper and we plan to actively maintain it for at least 5 years following the release date, after which the dataset will remain archived with its DOI.
+Appendix B Implementation Details
+We perform grid search to find the best hyperparameter settings for TemporalVLM on supervised dense video captioning on YouCook2 Zhou et al.
+(2018) and use them for all the remaining experiments.
+We present our hyperparameter settings in Tab.
+6.
 We sample $C=6$ clips from the video, with each consisting of $N^{c}_{f}=96$ frames.
@@
 Instead of sending timestamps from across the entire video, we send clip-wise timestamps along with their corresponding frames to the frame encoder and clip encoder respectively.
-Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q$ and overlap $o$.
-We then fuse the concatenated window-wise outputs of the Q-former via multiheaded self attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former.
+Inside the clip encoder, we divide the timestamp encoded input to the video Q-former into windows of size $q=32$ frames and overlap $o=16$ frames.
+We then fuse the concatenated window-wise outputs of the Q-former via multi-headed self-attention with 8 attention heads that have the same hidden size $N_{v}$ as the Q-Former.
 The output of the clip encoder contains a fused time-aware representation of each clip.
 We then concatenate the time-aware clip features according to their temporal order of appearance in the video.
-This sequence is of dimensions $(C*N^{c}_{f},~{}N_{V})$, with $N_{V}$ denoting the dimension of the video tokens.
-This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C*N^{c}_{f},~{}2*N_{V})$.
+This sequence is of dimensions $(C\times N^{c}_{f},~N_{V})$, with $N_{V}$ denoting the dimension of the video tokens.
+This sequence is passed to the BiLSTM module which uses information from both past and future states to output a global representation of dimensions $(C\times N^{c}_{f},~2\times N_{V})$.
 The BiLSTM module comprises of two hidden layers, with one hidden layer for the forward LSTM and another hidden layer for the backward LSTM.
 Projection layers are used to project the BiLSTM output into the LLM input space.
-The first layer projects the BiLSTM output to the dimensions of $(C*N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
+The first layer projects the BiLSTM output to the dimensions of $(C\times N^{c}_{f},N_{LLM})$ and then the second layer projects the output of the first layer to the sizes of $(N^{c}_{f},N_{LLM})$ which are the input dimensions required by the LLM.
 Hyperparameter Value
@@
 Number of frames sampled per clip $N^{c}_{f}$ 96
-Frame encoder patch size 14$\times$14
-Frame resolution 224$\times$224
+Frame encoder patch size $\times14$
+Frame resolution $\times224$
 Frame sampling type uniform
@@
 AdamW $\beta$ (0.9,0.999)
-Scheduler LinearWarmupCosineLR
 Q-Former window size $q$ 32
@@
 Number of layers in image encoder 12
-No of attention heads in fusion module 8
+Number of attention heads in fusion module 8
 Q-Former hidden size $N_{V}$ 768
@@
 Table 6: Hyperparameter settings.
-A.5 Comparison With Specialized Models
+Appendix C Comparisons with Specialized Models
 Tab.
-8 shows generalist models underperform specialized models such as Vid2Seq  [53] and QD-DETr [30] whose advantages come from task-specific training data and model designs.
-In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR  [30] leads to superior performance in highlight detection tasks.
-In contrast our TemporalVLM is trained on a simpler language modeling loss on lesser epochs.
-Generalist models such as ours exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
+7 shows that specialized models such as Vid2Seq Yang et al.
+(2023) and QD-DETR Moon et al.
+(2023), whose advantages come from task-specific training data and model designs, outperform generalist models such as TimeChat Ren et al.
+(2024) and our TemporalVLM model.
+In particular, the use of 4 loss functions for 200 epochs during training and the utilization of saliency tokens for saliency prediction in QD-DETR lead to superior performance in video highlight detection tasks.
+In contrast, our TemporalVLM model is trained on a simpler language modeling loss for a smaller number of epochs.
+Nevertheless, generalist models exhibit better generalization across zero-shot, multi-task, and multi-domain settings.
+Model Dense Captioning Highlight Detection Temporal Grounding
+YouCook2 QVHighlights Charades-STA
+SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
+Vid2Seq Yang et al. (2023) 7.9 47.1 27.3
+QD-DETR Moon et al. (2023) 38.9 62.4
+QD-DETR w/ Audio Moon et al. (2023) 39.0 62.9
+MMN Wang et al. (2022) 50.5 29.7
+VDI Luo et al. (2023a) 52.3 31.4
+Timechat Ren et al. (2024) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
+TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
+Table 7: Comparisons with specialized models. Best results are in bold, while second best ones are underlined.
+Appendix D Effects of Number of Training Epochs
+Tab.
+8 shows the performance of our model across different training epochs.
+Since our model uses a learnable fusion module in the clip encoder and a learnable BiLSTM module to aggregate clip-level features into a global representation, we observe the best results by training for 7 epochs.
+Tab.
+8 reports our results with 3 epochs, which are worse than ours with 7 epochs but still outperform those of TimeChat Ren et al.
+(2024) and LongVLM Weng et al.
+(2024).
+Model Epochs Dense Captioning Highlight Detection Temporal Grounding
+YouCook2 QVHighlights Charades-STA
+SODA_c CIDEr F1 mAP HIT@1 R@1 (IoU=0.5) R@1 (IoU=0.7)
+TimeChat Ren et al. (2024) 3 3.1 10.3 19.5 21.7 37.9 46.7 23.7
+LongVLM Weng et al. (2024) 3 2.3 8.1 16.9 16.0 22.5 27.2 11.9
+TemporalVLM 3 3.2 12.9 19.7 23.9 42.3 50.0 25.9
+TemporalVLM 7 3.4 13.2 20.0 25.1 43.0 54.4 29.0
+Table 8: Effects of number of training epochs. Best results are in bold, while second best ones are underlined.
+Appendix E General Video Understanding Results
+Following LongVLM Weng et al.
+(2024), we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT Maaz et al.
+(2023) in Tab.
+9.
+The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C).
+It is evident from Tab.
+9 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model.
+Model Data CI DO CU TU C
+Video-LLaMA Zhang et al. (2023) 10M 1.96 2.18 2.16 1.82 1.79
+Video-ChatGPT Ren et al. (2024) 100K 2.50 2.57 2.69 2.16 2.20
+Valley Luo et al. (2023b) 234K 2.43 2.13 2.86 2.04 2.45
+BT-Adapter Liu et al. (2024) 10M 2.68 2.69 3.27 2.34 2.46
+LongVLM Weng et al. (2024) 100K 2.76 2.86 3.34 2.39 3.11
+TemporalVLM 100K 2.88 2.91 3.45 2.50 3.16
+Table 9: General video understanding results. Best results are in bold, while second best ones are underlined.
+Appendix F Model Size Comparisons
+Tab.
+10 compares the sizes of our TemporalVLM model, TimeChat Ren et al.
+(2024), and LongVLM Weng et al.
+(2024) in terms of the number of learnable parameters (measured in millions).
+Our TemporalVLM model includes a learnable fusion module and a learnable BiLSTM module, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat and LongVLM respectively.
+Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat.
+In particular, TimeChat is trained on the complete TimeIT Ren et al.
+(2024) and Valley Luo et al.
+(2023b) datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download YT-Temporal Zellers et al.
+(2022) due to its large size and the restricted number of downloads.
+of the TimeIT and Valley datasets.
+Model Parameters
+TimeChat Ren et al. (2024) 241M
+LongVLM Weng et al. (2024) 244M
+TemporalVLM 255M
+Table 10: Model size comparisons.
+Appendix G Run Time Comparisons
+We compare the inference times (measured in seconds) of different models on three tasks, i.e., Dense Video Captioning (DVC) on YouCook2 Zhou et al.
+(2018), Video Highlight Detection (VHD) on QVHighlights Lei et al.
+(2021), and Temporal Video Grounding (TVG) on Charades Gao et al.
+(2017) in Tab.
+11.
+From the results, TimeChat Ren et al.
+(2024) is about 35% more efficient than LongVLM Weng et al.
+(2024) and TemporalVLM, since both LongVLM and TemporalVLM perform clip sampling and clip encoding.
 Model DVC VHD TVG
-SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
-Vid2Seq 7.9 47.1 27.3
-QD-DETR 38.9 62.4
-QD-DETR w/ Audio 39.0 62.9
-MMN 50.5 29.7
-VDI 52.3 31.4
-TemporalVLM 3.4 13.2 20.0 25.1 43.0 54.4 29.0
-Table 8: Comparisons with specialized models.
-A.6 Effects of Number of Training Epochs
+TimeChat Ren et al. (2024) 10.1s 5.6s 3.7s
+LongVLM Weng et al. (2024) 13.2s 7.3s 4.9s
+TemporalVLM 13.5s 7.6s 5.1s
+Table 11: Run time comparisons.
+Appendix H Effects of Number of Short-Term Clips
 Tab.
-9 shows the performance of our models across different training epochs.Since our model uses a learnable fusion layer in the clip encoder and a BiLSTM module to aggregate clip level features into a global representation, we observe the best results by training for 7 epochs.
-However, Tab.
-9 reports our results with 3 epochs, which underperform those with 7 epochs but still outperform those of TimeChat and LongVLM.
-Model DVC VHD TVG
-SODA_c CIDEr F1 mAP HIT@1 R@1 0.5 R@1 0.7
-LongVLM (3 epochs) 2.3 8.1 16.9 16.0 22.5 27.2 11.9
-TimeChat (3 epochs) 3.1 10.3 19.5 21.7 37.9 46.7 23.7
-TemporalVLM (3 epochs) 3.2 12.9 19.7 23.9 42.3 50.0 25.9
-TemporalVLM (7 epochs) 3.4 13.2 20.0 25.1 43.0 54.4 29.0
-Table 9: Effects of number of training epochs.
-A.7 Comparison on Model Sizes
-Tab.
-10 compares the sizes of our TemporalVLM model, TimeChat [36], and LongVLM [50] in terms of the number of learnable parameters.
-Our TemporalVLM model includes a trainable BiLSTM module and projection layers, leading to a 5% and 4% increase in the number of trainable parameters over TimeChat [36] and LongVLM [50] respectively.
-Nevertheless, our TemporalVLM model achieves the best performance across various temporal reasoning and fine-grained understanding tasks, despite using less training data than TimeChat [36].
-In particular, TimeChat [36] is trained on the complete TimeIT [36] and Valley [28] datasets, whereas our TemporalVLM model is trained on a subset [^2] We could not download the YT-Temporal [55] dataset on time due to its large size and the restricted number of downloads by YouTube.
-of the TimeIT [36] and Valley [28] datasets.
-A.8 Impacts of Number of Short-Term Clips
-Tab.
-11 presents the effects of the number of short-term clips on our TemporalVLM model.
+12 presents the effects of number of short-term clips on our TemporalVLM model.
 Due to memory limitations, we have experimented with three values for the number of short-term clips, namely 2, 4, and 6.
@@
 12 that the performance is improved with increasing the number of short-term clips, since our model is able to access more data from the input video.
-Model Number of Trainable Parameters
-TimeChat [36] 241 millions
-LongVLM [50] 244 millions
-TemporalVLM (Ours) 255 millions
-Table 10: Comparison on model sizes.
-# Clips SODA_c CIDEr F1
+Clips SODA_c CIDEr F1
 2 2.6 9.2 18.5
@@
 6 3.4 13.2 20.0
-Table 11: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
-A.9 General Video Understanding Results
-Method Data CI DO CU TU C
-Video-LLaMA [57] 10M 1.96 2.18 2.16 1.82 1.79
-Video-ChatGPT [29] 100k 2.50 2.57 2.69 2.16 2.20
-Valley [28] 234k 2.43 2.13 2.86 2.04 2.45
-BT-Adapter [25] 10M 2.68 2.69 3.27 2.34 2.46
-LongVLM [50] 100k 2.76 2.86 3.34 2.39 3.11
-TemporalVLM 100k 2.88 2.91 3.45 2.50 3.16
-Table 12: General video understanding results. Best results are in bold, while second best ones are underlined.
-Following LongVLM [50], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [29] in Tab.
-12.
-The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C).
-It is evident from Tab.
-12 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [50].
-A.10 Qualitative Results
+Table 12: Effects of number of short-term clips. Best results are in bold, while second best ones are underlined.
+Appendix I Results with a More Recent LLM
+The performance gain achieved by employing a more advanced LLM as the language decoder in our TemporalVLM model is illustrated in Fig.
+9.
+As LLama 3-8B is pretrained on a significantly larger and diverse dataset, it outperforms Llama 2-7B by a significant margin.
+The results demonstrate that as LLMs improve, the performance of VLMs, including our TemporalVLM model, will also improve alongside it.
+Figure 9: Results with a more recent LLM.
+Appendix J Qualitative Results
 In addition to Figs.
@@
 In particular, Figs.
-7, 8, and 9 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively.
+10, 11, and 12 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively.
 More importantly, qualitative results demonstrating the generalization abilities of TemporalVLM in the zero-shot setting are shown in Fig.
-10.
-Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understaning tasks.
+13.
+Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understanding tasks.
 Figure 10: Qualitative examples highlighting the dense video captioning capabilities of TemporalVLM in the supervised setting.
@@
 Blue denotes the ground truth and predicted timestamps.
-A.11 Discussions
-A.12 Limitations
-Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations.
-Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
-A.13 Societal Impacts
+Appendix K Limitations
+Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex temporal reasoning tasks and videos with extreme durations.
+In addition, our TemporalVLM model often has difficulty dealing with small objects, since it does not include an explicit object detector.
+Our future works will enhance TemporalVLM for tackling complex temporal reasoning tasks, videos with extreme durations, and dealing with small objects.
+Appendix L Societal Impacts
 Our time-sensitive video LLM could enable a variety of applications including optimizing and tracking industry assembly processes.